### v1

# CodeBrain: Bridging Decoupled Tokenizer and Multi-Scale Architecture for EEG Foundation Model

Abstract
Electroencephalography (EEG) provides real-time insights into brain activity and is widely used in neuroscience. However, variations in channel configurations, sequence lengths, and task objectives limit the transferability of traditional task-specific models. Although recent EEG foundation models (EFMs) aim to learn generalizable representations, they struggle with limited heterogeneous representation capacity and inefficiency in capturing multi-scale brain dependencies. To address these challenges, we propose CodeBrain, an efficient EFM structurally aligned with brain organization, trained in two stages. (1) We introduce a TFDual-Tokenizer that independently tokenizes heterogeneous temporal and frequency components, enabling a quadratic expansion of the discrete representation space. This also offers a degree of interpretability through cross-domain token analysis. (2) We propose the EEGSSM, which combines a structured global convolution architecture and a sliding window attention mechanism to jointly model sparse long-range and local dependencies. Unlike fully connected Transformer models, EEGSSM better reflects the brain’s small-world topology and efficiently captures EEG’s inherent multi-scale structure. EEGSSM is trained with a masked self-supervised learning objective to predict token indices obtained in TFDual-Tokenizer. Comprehensive experiments on 10 public EEG datasets demonstrate the generalizability of CodeBrain with linear probing. By offering biologically informed and interpretable EEG modeling, CodeBrain lays the foundation for future neuroscience research. Both code and pretraining weights will be released in the future version.
1Saw Swee Hock School of Public Health, National University of Singapore, Singapore
2Institute of Data Science, National University of Singapore, Singapore
3College of Computing and Data Science, Nanyang Technological University, Singapore
4Beijing Key Laboratory of Brainnetome and Brain-Computer Interface, Institute of Automation, Chinese Academy of Sciences, Beijing, China
5Brainnetome Center, Institute of Automation, Chinese Academy of Sciences, Beijing, China
1 Introduction
Electroencephalography (EEG) data represent the brain’s electrical activity, captured by placing electrodes on the scalp niedermeyer2005electroencephalography . With its high temporal resolution, EEG provides real-time insights into brain dynamics, supporting neuroscience and cognitive research da2013eeg ; binnie2003cognitive . To enable automated analysis, researchers have developed various task-specific models, including sleep staging phan2021xsleepnet ; lee2025explainable ; ma2024st , emotion recognition jia2020sst ; guo2024eeg ; wang2024dmmr , motor imagery wang2024st ; ari2024nf ; li2020learning , and other applications guerra2024exploring ; hu2024exploring . However, these models are usually trained from scratch and tend to capture task-specific patterns, limiting their applicability beyond the original task. Moreover, variations in channel configurations and input lengths across EEG tasks further hinder knowledge transfer. To tackle these issues, EEG foundation models (EFMs) are developed to learn universal representations for diverse downstream tasks zhou2025brain .
Inspired by masked self-supervised pretraining in natural language processing van2017neural ; devlin2019bert , current EFMs commonly adopt patch-wise representation learning: EEG signals are divided into fixed-length patches, encoded into latent representations, and trained to reconstruct the masked portions. This form of pretraining offers flexibility in downstream applications with varying channel configurations or input lengths by adjusting patch arrangement and number. Among reconstruction-based methods, some directly predict the masked raw signals wang2024eegpt ; wang2024cbramod , a task that is challenging due to EEG’s inherent noise and variability. Others adopt vector quantization to abstract away signal-level fluctuations jianglarge ; pradeepkumar2025single , and follow a two-stage pipeline: 1) Learn a tokenizer with a codebook to quantize EEG patches into discrete tokens, and 2) pretrain the EFM to predict the token indices of masked patches. By acting as a noise-robust latent space, the codebook enables the model to learn more expressive representations. Despite these advances, existing EFMs still face challenges.
Failing to Decouple Heterogeneous EEG Representations. Recent EFMs adopt vector quantization with a single codebook for noise-robust representation jianglarge ; pradeepkumar2025single , following image-based VQ-VAE architecture van2017neural , where spatially entangled and homogeneous visual features make a single codebook sufficient zheng2023online ; mentzerfinite . However, EEG signals exhibit heterogeneous structures: temporal and frequency components reflect distinct aspects of brain activity miwakeichi2004decomposing . Applying a unified codebook to such heterogeneous data can blur domain-specific characteristics, potentially conflating localized temporal events with global frequency patterns liu2024vector and limiting the representation capacity.
Struggling with Modeling Global Brain Dependencies. EEG signals exhibit sparse global dependencies and strong local correlations, reflecting the brain’s small-world topology bullmore2009complex ; bassett2006small ; he2009uncovering . Efficient modeling of such structure requires capturing relationships in a scalable way. However, most EFMs yang2023biot ; jianglarge ; mohammadi2024eeg2rep ; wang2024eegpt ; wang2024cbramod adopt Transformer architectures vaswani2017attention with fully connected self-attention. This over-connected design is misaligned with the brain’s sparse structure and struggles to efficiently capture global dependencies due to its quadratic complexity with sequence length tay2021long ; hong2025eegm2 ; tegon2025femba .
Neglecting Local Dependencies within EEG Patches. EEG signals exhibit rich local waveform structures over short temporal windows, reflecting crucial transient neural events tatum2021handbook . However, most existing EFMs represent each EEG patch as a single token and apply attention mechanisms only at the patch level wang2024cbramod ; jianglarge , thereby ignoring important local dependencies within patches.
Figure 1: Overview of our contributions. (a) Decoupled vector quantization independently tokenizes temporal and frequency components to preserve heterogeneous EEG structures and enhance representation capacity. (b) State Space Model efficiently captures sparse global dependencies across patches. (c) Sliding Window Attention models fine-grained local dependencies within patches.
To address the above challenges, we propose CodeBrain, an efficient EEG foundation model that integrates a decoupled time–frequency tokenizer and a multi-scale architecture. CodeBrain is trained in two stages. In the first stage, we introduce TFDual-Tokenizer (Fig. 1(a)), which independently tokenizes heterogeneous temporal and frequency EEG components. To stabilize training, we apply a contrastive learning objective to optimize the temporal codebook. In the second stage, we propose EEGSSM, a framework for learning multi-scale EEG representations by masked self-supervised learning. EEGSSM predicts token indices obtained from the TFDual-Tokenizer for masked patches. To address the inefficiency of over-connected attention and better align with the brain’s small-world structure, we introduce Structured Global Convolution (SGConv) (Fig. 1(b)) as the backbone. Inspired by Structured State Space Models gu2021efficiently ; smith2023simplified ; li2022makes ; gu2022parameterization , SGConv performs efficient global convolution over long EEG sequences, naturally capturing long-range dependencies and local clustering patterns characteristic of brain networks. Complementarily, we integrate a Sliding Window Attention (SWA) mechanism (Fig. 1(c)), adapted from long-sequence modeling beltagy2020longformer , to capture local intra-patch dependencies. Together, SGConv and SWA enable EEGSSM to perform efficient and scalable multi-scale EEG representation learning. Our detailed contributions can be summarized as follows:
  • Decoupled Heterogeneous Tokenizer. We propose the TFDual-Tokenizer, which independently encodes temporal and frequency components of EEG signals, leading to a quadratic expansion of the discrete representation space. It also provides interpretability, supported by visualization analyses in Section 4.5. To the best of our knowledge, this is the first vector quantization framework that decouples heterogeneous domains in EEG foundation models.
  • Efficient Multi-Scale Architecture. We design an efficient architecture combining a structured global convolution architecture backbone for global inter-patch dependencies and sliding window attention mechanism for local intra-patch dependencies. This design aligns with the brain’s small-world topology, enabling scalable multi-scale representation learning.
  • Comprehensive Validation. We evaluate CodeBrain on 8 diverse downstream brain-computer interface tasks using 10 public EEG datasets. Our CodeBrain model demonstrates strong generalization and state-of-the-art performance with linear probing.
2 Related Work
EEG Foundation Models. Inspired by foundation models in vision and language wang2023internimage ; achiam2023gpt , EEG research is shifting from task-specific models jia2020graphsleepnet ; wang2024caresleepnet ; chen2025long to EFMs for learning expressive representations. Current EFMs can be divided into two categories. 1) Contrastive learning-based (CL) chen2025large : BENDER kostas2021bendr first showed the ability of CL for EEG representations. Then, the Brant series zhang2023brant ; yuan2024brant ; zhang2024brant enables joint representation learning across physiological signals using CL. 2) Reconstruction-based: BIOT yang2023biot pioneers cross-modal pretraining for biosignals, including EEG. Subsequent models focus specifically on EEG, learning representations by predicting masked discrete tokens jianglarge ; jiangneurolm or reconstructing raw signals wang2024eegpt ; mohammadi2024eeg2rep ; wang2024cbramod . Most existing EFMs adopt Transformer architecture, which is suboptimal for EEG due to poor handling of sparse dependencies, quadratic complexity, and ignoring local dependencies by treating each patch as a token.
EEG Tokenization. Tokenization has been key in NLP for generating generalizable and interpretable input representations sennrich2016neural ; kudo2018sentencepiece . Inspired by this, early EFMs used patch-based continuous tokenization yang2023biot ; yuan2024brant to handle EEG noise and variability, but without quantization, leading to unbounded and less interpretable representations. LaBraM jianglarge introduced vector quantization to learn discrete EEG tokens, following the VQ-VAE design from vision tasks van2017neural . However, this direct transfer overlooks EEG’s heterogeneous structure, limiting representation capacity. Moreover, the tokenizer of LaBraM is trained only on frequency-domain reconstruction due to convergence issues with raw signals reconstruction. Later efforts pradeepkumar2025single ; jiangneurolm all relied on a single codebook to encode mixed domain features, limiting both the representation space and interpretability.
State Space Model Recent works have focused on enhancing classic State-Space Models (SSM) to more efficiently model sequential data using deep learning. For example, Syama et al. rangapuram2018deep used recurrent neural networks to learn parameters in SSM. However, the most significant progress came from Gu et al. gu2021efficiently with the introduction of the structured state space model (S4), which reduces the computational complexity of SSM in modeling long sequences using a special state transition matrix gu2020hippo . These low-rank and normal matrices enable SSM to compute global convolution kernels efficiently through fast Fourier transform across the entire sequence. Subsequently, some works have further improved the shortcomings of S4 in areas such as model architecture smith2022simplified and convolution raghu2023sequential , and have started applying it to tasks such as natural language processing dao2022hungry and time series analysis zhou2023deep . Recently, some works have also been applied to the EEG data, Tran et al. tran2024eeg leveraging SSMs to detect dementia. They extract temporal information from EEG signals through the Mamba architecture and combine it with frequency domain features to better manage the complexity of multivariate EEG. In the work of Gui et al. gui2024eegmambabidirectionalstatespace , SSM has also become the backbone network of the EEG foundation model. This further highlights the fast reasoning speed and efficient memory usage of the SSM model when processing EEG signals.
3 Methodology
3.1 Preliminaries
Decoupled Tokenization. Given a normalized EEG patch $X\in\mathbb{R}^{L}$, where $L$ is the patch length, our goal is to learn a tokenizer that maps $X$ to discrete tokens using two separate codebooks: one for the temporal domain and one for the frequency domain. Specifically, let $V^{(t)}\in\mathbb{R}^{K\times D}$ and $V^{(f)}\in\mathbb{R}^{K\times D}$ denote the temporal and frequency codebooks, respectively, where $K$ is the vocabulary size and $D$ is the embedding dimension of each token. The tokenizer function is defined as: $z^{(t)},z^{(f)}=f_{\text{tokenizer}}(X)$, where $z^{(t)},z^{(f)}\in\{1,\dots,K\}$ are the indices of the selected codes from the temporal and frequency codebooks.
EEG Representation Learning. Given unlabeled EEG sequences $\mathcal{X}=\{X_{i}\}_{i=1}^{N}$, where each sequence $X_{i}\in\mathbb{R}^{C\times T\times N}$ consists of $C$ channels, a sampling rate $T$, and a duration of $N$ seconds, the goal is to train an encoder $f_{\text{enc}}:\mathbb{R}^{C\times T\times N}\rightarrow\mathbb{R}^{D}$ that produces latent representations $Z_{i}=f_{\text{enc}}(X_{i})$.
3.2 Model Architecture
We introduce CodeBrain, a two-stage pretraining framework designed to address the two problem statements described in Section 3.1. The model architecture of CodeBrain is illustrated in Fig. 2. In the first stage, we propose TFDual-Tokenizer to decouple temporal and frequency representations of EEG signals, converting the raw signals into noise-robust tokens with potential interpretability. In the second stage, we introduce the EEGSSM framework for EEG representation learning.
Figure 2: Overview of the CodeBrain framework. Left: TFDual-Tokenizer learns to discretize EEG signals into temporal and frequency tokens using two separate codebooks, by reconstructing both the temporal waveforms and the frequency-domain magnitude and phase. Right: EEGSSM learns representations by predicting the discrete tokens of masked patches generated by TFDual-Tokenizer.
3.3 TFDual-Tokenizer Pretraining
Our TFDual-Tokenizer includes a shared neural encoder, a dual tokenizer with separate codebooks, and two decoders. The neural encoder extracts joint time-frequency embeddings from EEG patches, which are then discretized into temporal and frequency tokens by the dual tokenizer. Each token stream is reconstructed by a decoder to supervise codebook learning in its respective domain.
Neural Encoder For each patched and normalized EEG input $X\in\mathbb{R}^{C\times T\times N}$, we apply the Discrete Fourier Transform (DFT) cooley1965algorithm to obtain its frequency representation:
$X[k]=\sum_{n=0}^{N-1}x_{n}\cdot e^{-j\frac{2\pi}{N}kn},\quad k=0,1,\dots,N-1,$ (1)
where $X[k]$ denotes the $k$-th frequency component, and $n$ denotes the number of patches. $X[k]$ is then combined with the raw time-domain signal and fed into the TFConv module, where both are processed in parallel through stacks of convolutional, batch normalization, and ReLU layers. The outputs $e_{t}$ and frequency representation $X[K]$ are concatenated to form a time-frequency embedding $e_{p}=Concat\{e_{t},X[K]\}$. To obtain patch representation $\tilde{e}_{n}$, we then add a positional embedding $e_{\text{pos}}$ and feed them into TF-Encoder, a Transformer encoder:
$\tilde{e}_{n}=TransformerEncoder(e_{p}+e_{pos}).$ (2)
Dual Tokenizer We use two separate tokenizers with distinct codebooks for the temporal and frequency domains, denoted as $v^{(t)}_{j},v^{(f)}_{j}\in\mathbb{R}^{D}$, where $j=1,\dots,K$. Given the patch representation $\tilde{e}_{i}$ from the neural encoder, each tokenizer independently selects the nearest code from its codebook:
$\hat{z}^{(t)}_{i}=\arg\min_{j}\|\tilde{e}_{i}-v^{(t)}_{j}\|^{2},\quad\hat{z}^{(f)}_{i}=\arg\min_{j}\|\tilde{e}_{i}-v^{(f)}_{j}\|^{2}.$ (3)
Frequency Codebook Training To train the frequency codebook, we reconstruct two fundamental features—amplitude and phase—from the code embeddings. For each EEG patch, we apply the DFT (Eq. (1)) to obtain the complex frequency representation: $X[k]=\text{Re}\{X[k]\}+j\cdot\text{Im}\{X[k]\}$
where $\text{Re}\{X[k]\}$ and $\text{Im}\{X[k]\}$ are the real part and imaginary part respectively, then the amplitude and phase can be calculated as:
$A_{n}=\sqrt{\text{Re}(X[k])^{2}+\text{Im}(X[k])^{2}},\quad\phi_{n}=\arctan 2\left(\text{Im}(X[k]),\text{Re}(X[k])\right).$ (4)
We use z-score normalization to ensure stable training. The code embedding $v^{(f)}_{\hat{z}_{n}}$, retrieved from the frequency codebook, is passed through the F-Decoder, which consists of a Transformer encoder followed by two linear layers:
$y_{n}^{A}=TF^{(f)}(MLP(\tilde{e}_{n})),\quad y_{n}^{P}=TF^{(f)}(MLP(\tilde{e}_{n})).$ (5)
where $y_{n}^{A}$ and $y_{n}^{P}$ are the predicted amplitude and phase, respectively. $TF^{(f)}$ denotes a Transformer encoder. The frequency codebook’s training objective frequency codebook is the mean squared error (MSE) loss:
$\mathcal{L}_{n}^{f}=||y_{n}^{A}-A_{n}||^{2}_{2}+||y_{n}^{P}-\phi_{n}||^{2}_{2}.$ (6)
Temporal Codebook Training Direct reconstruction of temporal features might lead to non-convergence jianglarge . To address this, we combine contrastive loss with reconstruction loss to train a temporal codebook. Inspired by studies on physiological signals kiyasseh2021clocs , we assume temporal dependencies exist between different EEG patches, especially within the same channel. For an EEG sample $X\in\mathbb{R}^{C\times T\times N}$, we split it into two halves of length $N/2$. We use TF-Encoder in Eq. (2) to obtain the representations of these two segments $X_{h1},X_{h2}\in\mathbb{R}^{C\times T\times N/2}$ separately:
$h_{n}=TransformerEncoder(X_{{n}}),X_{n}\in\mathbb{R}^{C\times(T\times N/2)}.$ (7)
We encourage the latent representations of different segments within a single sample $h_{n}$ and $\tilde{h}_{n}$ to be similar, while making the latent representations of the same segment across different samples $h_{n}$ and $h_{k}$ as distinct as possible. We use SimCLR loss chen2020simple to train the temporal codebook:
$\mathcal{L}_{CL}=-\log\frac{\exp\left(\mathrm{sim}({h}_{n},\tilde{h}_{n})/\tau\right)}{\sum\limits_{k=1}^{2N}\mathbb{1}_{[k\neq i]}\exp\left(\mathrm{sim}({h}_{n},{h}_{k})/\tau\right)}\quad,$ (8)
where $\mathrm{sim}(\mathbf{h}_{i},\mathbf{h}_{j})$ is the cosine similarity, $\tau>0$ is the temperature parameter, and $\mathbb{1}_{[k\neq i]}$ is an indicator function used to exclude itself. We introduce a T-Decoder to reconstruct raw signals from the temporal code embedding $\tilde{e}_{n}$, which consists of a Transformer encoder $TF^{(t)}$ followed by a linear projection. Let $y_{t}$ be the Transformer output from Eq. (2); the overall training objective is:
$\mathcal{L}_{t}=\mathcal{L}_{CL}+||y_{t}-x_{t}||^{2}_{2},\quad y_{t}=TF^{(t)}(MLP(\tilde{e}_{n})).$ (9)
Finally, the training objective for the TFDual-Tokenizer is defined as:
$\mathcal{L}_{codebook}=\sum_{x\in\mathcal{D}}\sum_{i}\underbrace{||y_{n}^{A}-A_{n}||^{2}_{2}+||y_{n}^{P}-\phi_{n}||^{2}_{2}}_{\text{reconstruction loss}}+\underbrace{\|sg(\tilde{e}_{n})-v_{j}^{(t)}\|^{2}+\|sg(\tilde{e}_{n})-v_{j}^{(f)}\|^{2}}_{\text{codebook loss}}+\mathcal{L}_{t},$ (10)
where $sg(\cdot)$ denotes stop-gradient, avoiding the update of encoder parameters.
3.4 EEGSSM Pretraining
In this stage, we introduce a novel convolutional structured state space model framework, called EEGSSM, composed of multiple EEGSSM blocks. First, raw EEG signals are projected into a higher-dimensional embedding space via a 1D convolutional layer. The resulting features are processed by EEGSSM blocks. Another 1D convolutional layer then maps the output back to the token space for reconstructing the indices of masked tokens produced by the TFDual-Tokenizer.
EEGSSM Block Our EEGSSM block is composed of several blocks, which are integrated together through a residual connection mechanism. An EEGSSM block consists of an RMS Norm layer, SGConv layer, SWA layer, and a gating component. Root Mean Square Layer Normalization (RMSNorm) zhang2019rootmeansquarelayer can be considered an improved version of LayerNorm, normalizing by using the square root of the mean to reduce the impact of noise. Afterward, we feed the intermediate variables into the SGConv layer to obtain a global receptive field through convolution SSM.
SGConv Layer. SGConv is a structured SSM model using convolution architecture (See supplemental material), and its convolution structure can be represented as a FFT formula:
$y=F^{-1}_{N}D_{k}F_{N}u,D_{k}=\text{diag}(\overline{K}F_{N}),$ (11)
where $F_{N}$ denotes the DFT matrix of size $N$. This FFT convolution has a computational complexity of $O(nlog(n))$. As a type of convolutional SSM, SGConv improves the convolution kernel $\overline{K}$ in Eq. (11) by introducing two features: sparse parameterization and kernel decay, making SGConv easier and more efficient to compute compared to the traditional S4 kernel. Let $L$ be the length of the input sequence. The convolution kernel $\overline{K}$ of SGConv is composed of several sub-kernels. Assuming the size of the first sub-kernel is $d$, with parameters $w_{i}\in\mathbb{R}^{d\times d}$, then the number of sub-kernels can be expressed as $N=\log_{2}(\frac{L}{d})+1$. The convolution kernel $\overline{K}$ in Eq. (11) can thus be initialized as:
$\overline{K}=\frac{1}{Z}[k_{0},k_{1},...,k_{N-1}],k_{i}=\alpha^{i}\mathrm{Upsample}_{2^{max[i-1,0]}d}(w_{i}),$ (12)
where $\alpha$ denotes the decay coefficient, usually chosen to be 1/2, inducing the decaying structure, and $\text{Upsample}(x)$ denotes upsample $x$ to length $l$. We also introduce the normalization constant $Z$ to ensure that the convolution operation does not change the scale of the input.
Sliding Window Attention Layer. We included a sliding window attention (SWA) layer to capture fine-grained local temporal dependencies. We apply a small fixed-length window, allowing the model to directly access the content within the context of the window through an attention mechanism by sliding it across the entire sequence, thereby addressing previous models’ neglect of intra-patch temporal information. Furthermore, SWA maintains linear computational complexity to ensure training speed remains largely unaffected.
At the end of the EEGSSM block, we use a gating mechanism to control the output of the block. We employ a gated unit similar to Wavenet obeid2016temple , which can suppress useless or irrelevant features and help stabilize training in deep networks. We concatenate the global features output $y_{sg}$ by the SGConv layer with the local features output $y_{swa}$ by SWA and feed them into a gated unit:
$z=tanh(W_{f}\times Concat(y_{sg},y_{swa}))\odot\sigma(W_{g}\times Concat(y_{sg},y_{swa})),$ (13)
$y_{1}=\mathrm{Conv}(z),y_{2}=\mathrm{Conv}(z)$ (14)
where $tanh(\cdot)$ and $\sigma(\cdot)$ are the tanh function and sigmoid function, $\odot$ denotes an element-wise multiplication operator, $W_{f}$ and $W_{g}$ are learnable convolution filters. $y_{1}$ becomes the input to the next block, while $y_{2}$ will be aggregated to the output of SSM blocks through a skip connection.
Pre-Training Objective To help the EEGSSM model learn general EEG representations, we use a Masked Autoencoder (MAE) for self-supervised pre-training. For a patched sample $X=\{x_{i}\mid i\in[1,2,\ldots,C]\}$, we randomly generate a mask =$\mathcal{M}=\{m_{i}\mid i\in[1,2,\ldots,C]\}$ from a Bernoulli distribution of $r$ proportion, where $m_{ij}\in\{0,1\}$. We reconstruct the token indices of masked EEG patches from the TFDual-Tokenizer by cross-entropy loss. Let $y_{i}$ denote the output of the EEGSSM block, the probability that the EEG signal matches the corresponding token $v_{i}$ in the codebooks:
$p(v_{i}|x_{i})=softmax(Conv1D(y_{i})).$ (15)
Suppose the size of the pre-training set is $N$, the final cross-entropy loss is:
$\mathcal{L}_{p}=-\sum_{j=0}^{N}\sum_{n\in\{m_{i}=1\}}^{C}p(v_{nj}|x_{nj}).$ (16)
4 Experiments
4.1 Datasets
Pre-Training We pretrain CodeBrain on the TUH EEG Corpus (v2.0.1, modified in March 2025) obeid2016temple , the largest publicly available EEG dataset to date. Data processing follows a standardized pipeline: recordings shorter than 5 minutes are excluded, and the first and last minute of each segment is removed. We retain 19 commonly used EEG channels (C3, C4, Cz, F3, F4, Fp1, Fp2, F7, F8, Fz, O1, O2, P3, P4, Pz, T3, T4, T5, T6), selected based on the international 10–20 system for electrode placement hh1958ten , apply band-pass filtering (0.3–75 Hz), and notch filtering at 60 Hz to remove noise. The data is resampled to 200 Hz and divided into 30-second non-overlapping segments. Segments with absolute amplitudes over $100\leavevmode\nobreak\ \mu\mathrm{V}$ are filtered out. To normalize the signals, each value is divided by 100. Each segment is split into 1-second windows, resulting in 570 EEG patches per sample. After preprocessing, 1,109,545 samples (about 9,246 hours) are retained for pretraining.
Downstream Tasks. We evaluate CodeBrain on 8 downstream tasks across 10 publicly available EEG datasets, covering a diverse range of applications. Detailed dataset configurations are in Table 1. We perform cross-subject or cross-session splits with strict separation between training, validation, and test sets. For FACED chen2023large , we use 80 subjects for training, 20 for validation, and the remaining 23 for testing. In SEED-V liu2021comparing , we perform a leave-one-session-out manner. ISRUC_S3 khalighi2016isruc consists of 10 subjects, for which we apply an 8:1:1 cross-subject split. MentalArithmetic consists of 36 subjects, and we use a 7:1:1 cross-subject split. BCIC2020-T3 jeong20222020 follows the official competition protocol. In CHB-MIT shoeb2009application , we use recordings from 19 subjects for training, and from 2 subjects each for validation and testing. Additional dataset details are provided in the supplementary material.
Table 1: Summary of Downstream Tasks and Associated EEG Datasets.
Downstream Tasks Datasets Shape #Samples Class
Emotion Recognition FACED chen2023large 32 channels * 10s 10,332 9-class
SEED-V liu2021comparing 62 channels * 1s 117,744 5-class
Sleep Staging ISRUC_S1 khalighi2016isruc 6 channels * 30s 86,320 5-class
ISRUC_S3 khalighi2016isruc 6 channels * 30s 8,500 5-class
Imagined Speech Classification BCIC 2020-T3 jeong20222020 64 channels * 3s 6,000 5-class
Mental Stress Detection Mental Arithmetic mumtaz2016eeg 20 channels * 5s 1,707 2-class
Seizure Detection CHB-MIT shoeb2009application 16 channels * 10s 326,993 2-class
Motor Imagery Classification SHU-MI goldberger2000physiobank 32 channels * 4s 11,988 2-class
Event Type Classification TUEV obeid2016temple 16 channels * 5s 112,491 6-class
Abnormal Detection TUAB obeid2016temple 16 channels * 10s 409,455 2-class
4.2 Experiment Settings
Two-Stage Training Setup. All experiments are conducted on NVIDIA 40GB A100 GPUs. In Stage 1, the TFDual-Tokenizer is trained with temporal and frequency codebooks of 4096 codes (32 dimensions) for 20 epochs, using a batch size of 256 and a learning rate of 1e-4, across six A100 GPUs for approximately ten hours. In Stage 2, an 8-layer EEGSSM backbone with a masking ratio of 0.5 is trained for 10 epochs, using a batch size of 256 on two A100 GPUs for about one day.
Fintuning Strategy. We adopt a linear probing setup to evaluate the quality of pretrained representations. All downstream task datasets are resampled to 200 Hz to match the pretraining configuration. A three-layer MLP is applied. The first layer aggregates information across all EEG channels. The second layer compresses the entire $x$-second sequence (depending on task configuration) into a 200-dimensional vector. The final layer maps this representation to the target class. Each hidden layer uses ELU activation and dropout to improve generalization.
Baselines. We compare our model with five publicly available EFMs that have released pretrained weights, representing different pretraining strategies. BENDR kostas2021bendr adopts contrastive learning to learn general EEG representations. BIOT yang2023biot uses patch-based continuous tokenization. LaBraM introduces discrete neural tokens through vector quantization jianglarge . EEGPT wang2024eegpt and CBraMod wang2024cbramod both rely on masked reconstruction of raw EEG signals.
Evaluation Metric. For multi-class classification, we report Cohen’s Kappa, Weighted F1 score, and Balanced Accuracy, with Kappa based on validation performance for testing. For binary classification, we use Area Under the ROC Curve (AUROC), Area Under the Precision-Recall Curve (AUC-PR), and Balanced Accuracy, with AUROC based on validation performance for testing. All experiments are repeated with five random seeds, and we report the mean and standard deviation.
4.3 Comparison with Baselines
We ensure consistent data splits across all baselines. Results are reported on six representative downstream datasets, with additional results provided in the supplementary material. As shown in Table 2, CodeBrain achieves consistent performance gains compared to baselines. For multi-class classification, it achieves the largest gain of +0.0911 in Cohen’s Kappa (21.6%), +0.0718 in Weighted F1 score (13.3%), +0.0728 in Balanced Acc (13.5%) on BCIC 2020-T3 over the strongest baseline wang2024cbramod . For binary classification, it achieves the largest gain of +0.0802 in AUROC (10.1%), +0.0910 in AUC-PR (14.5%) and +0.0258 in Balanced Acc (3.6%) on Mental Arithmetic over the strongest baseline wang2024cbramod . These results demonstrate the superior generalizability of CodeBrain.
Table 2: Comparison results of different methods on downstream tasks.
Methods FACED (9-Class) SEED-V (5-Class)
Kappa W-F1 Bal. Acc Kappa W-F1 Bal. Acc
BENDRkostas2021bendr 0.4716 ± 0.0095 0.5340 ± 0.0086 0.5320 ± 0.0083 0.0335 ± 0.0062 0.2026 ± 0.0330 0.2231 ± 0.0059
BIOTyang2023biot 0.4476 ± 0.0254 0.5136 ± 0.0112 0.5118 ± 0.0118 0.2261 ± 0.0262 0.3856 ± 0.0203 0.3837 ± 0.0187
LaBraMjianglarge 0.4698 ± 0.0102 0.5288 ± 0.0188 0.5273 ± 0.0107 0.2386 ± 0.0209 0.3974 ± 0.0111 0.3976 ± 0.0138
EEGPTwang2024eegpt 0.4639 ± 0.0023 0.3924 ± 0.0017 0.4607 ± 0.0014 0.1323 ± 0.0062 0.3090 ± 0.0052 0.3061 ± 0.0044
CBraModwang2024cbramod 0.5041 ± 0.0122 0.5618 ± 0.0093 0.5509 ± 0.0089 0.2569 ± 0.0143 0.4101 ± 0.0108 0.4091 ± 0.0097
CodeBrain 0.5270 ± 0.0039 0.5849 ± 0.0033 0.5816 ± 0.0028 0.2735 ± 0.0032 0.4235 ± 0.0022 0.4137 ± 0.0023
Methods ISRUC_S3 (5-Class) BCIC 2020-T3 (5-Class)
Kappa W-F1 Bal. Acc Kappa W-F1 Bal. Acc
BENDRkostas2021bendr 0.5995 ± 0.0151 0.6789 ± 0.0142 0.6352 ± 0.0095 0.0607 ± 0.0093 0.2379 ± 0.0165 0.2485 ± 0.0075
BIOTyang2023biot 0.7168 ± 0.0119 0.7834 ± 0.0096 0.7598 ± 0.0109 0.3650 ± 0.0176 0.4917 ± 0.0079 0.4920 ± 0.0086
LaBraMjianglarge 0.7194 ± 0.0162 0.7843 ± 0.0189 0.7617 ± 0.0122 0.3800 ± 0.0242 0.5054 ± 0.0205 0.5060 ± 0.0155
EEGPTwang2024eegpt 0.6160± 0.0856 0.6375 ± 0.0632 0.6650 ± 0.0311 0.0567 ± 0.0164 0.2441 ± 0.0105 0.2453 ± 0.0131
CBraModwang2024cbramod 0.7407 ± 0.0251 0.8056 ± 0.0219 0.7844 ± 0.0126 0.4216 ± 0.0163 0.5383 ± 0.0096 0.5373 ± 0.0108
CodeBrain 0.7671 ± 0.0091 0.8202 ± 0.0071 0.7856 ± 0.0031 0.5127 ± 0.0065 0.6101 ± 0.0053 0.6101 ± 0.0052
Methods Mental Arithmetic (2-Class) CHB_MIT (2-Class)
AUROC AUC-PR Bal. Acc AUROC AUC-PR Bal. Acc
BENDRkostas2021bendr 0.6248 ± 0.0765 0.3661 ± 0.0672 0.5681 ± 0.0448 0.8632 ± 0.0526 0.3071 ± 0.1240 0.5609 ± 0.0432
BIOTyang2023biot 0.7536 ± 0.0144 0.6004 ± 0.0195 0.6875 ± 0.0186 0.8761 ± 0.0284 0.3277 ± 0.0460 0.7068 ± 0.0457
LaBraMjianglarge 0.7721 ± 0.0093 0.5999 ± 0.0155 0.6909 ± 0.0125 0.8679 ± 0.0199 0.3287 ± 0.0402 0.7075 ± 0.0358
EEGPTwang2024eegpt 0.7162 ± 0.0171 0.5081 ± 0.0275 0.5597 ± 0.0171 0.8892 ± 0.0066 0.3073 ± 0.0641 0.5481 ± 0.0151
CBraModwang2024cbramod 0.7905 ± 0.0073 0.6267 ± 0.0099 0.7256 ± 0.0132 0.8892 ± 0.0154 0.3689 ± 0.0382 0.7398 ± 0.0284
CodeBrain 0.8707 ± 0.0209 0.7177 ± 0.0421 0.7514 ± 0.0203 0.8961 ± 0.0174 0.4377 ± 0.0288 0.7273 ± 0.0240
4.4 Ablation Study
We conduct ablation studies on three datasets using the same five seeds as in the main experiments to evaluate the key components in CodeBrain, as shown in Table 3. Additional results on mask ratio and architecture efficiency are in the supplementary material. Below is the detailed analysis:
(1) Tokenizer configuration: we compare the proposed TFDual-Tokenizer (Dual) with variants using only one domain codebook (Temporal or Frequency), or a shared codebook jointly reconstructing both domains (Mixed). Across all datasets, the Dual codebook consistently yields superior performance, suggesting that maintaining separate temporal and frequency codebooks captures complementary information more effectively than merging them.
(2) Sliding Window Attention (SWA): Including the SWA module consistently improves performance, confirming its regularization effect in capturing local dependencies. Notably, on the SEED-V dataset with 1s samples, SWA yields the largest gain, confirming its effectiveness for short-duration inputs.
(3) Contrastive learning (CL): We evaluate the impact of contrastive learning in TFDual-Tokenizer pretraining. It also leads to consistent gains, indicating a better capture of temporal patterns.
(4) Scaling Laws: Prior work wang2024cbramod ; jianglarge has explored scaling with 1 to 1000 hours of EEG data for pretraining and model sizes up to 5M parameters. We extend this by analyzing performance using 1k to 9k hours of pretraining data and larger models. As shown in Figure 3, Kappa, our primary evaluation metric, consistently improves with more data and parameters, suggesting that further scaling may lead to continued performance gains.
Table 3: The results of ablation studies for tokenizer configurations and module components.
Dataset Codebook SWA CL Cohen’s Kappa Weighted F1 Balanced Accuracy
FACED 9-Class Dual 0.5270 $\pm$ 0.0039 0.5849 $\pm$ 0.0033 0.5816 $\pm$ 0.0028
Temporal 0.4618 $\pm$ 0.0072 0.5277 $\pm$ 0.0067 0.5217 $\pm$ 0.0056
Frequency 0.5006 $\pm$ 0.0224 0.5607 $\pm$ 0.0201 0.5580 $\pm$ 0.0187
Mixed 0.4676 $\pm$ 0.0061 0.5319 $\pm$ 0.0052 0.5281 $\pm$ 0.0049
Dual 0.5222 $\pm$ 0.0082 0.5811 $\pm$ 0.0084 0.5765 $\pm$ 0.0074
Dual 0.5192 $\pm$ 0.0092 0.5792 $\pm$ 0.0093 0.5736 $\pm$ 0.0075
SEED-V 5-Class Dual 0.2735 $\pm$ 0.0032 0.4235 $\pm$ 0.0022 0.4137 $\pm$ 0.0023
Temporal 0.2633 $\pm$ 0.0116 0.4152 $\pm$ 0.0092 0.4068 $\pm$ 0.0074
Frequency 0.2665 $\pm$ 0.0208 0.4186 $\pm$ 0.0177 0.4098 $\pm$ 0.0147
Mixed 0.2708 $\pm$ 0.0047 0.4214 $\pm$ 0.0044 0.4124 $\pm$ 0.0032
Dual 0.2589 $\pm$ 0.0065 0.4129 $\pm$ 0.0056 0.4029 $\pm$ 0.0042
Dual 0.2561 $\pm$ 0.0051 0.4106 $\pm$ 0.0042 0.4019 $\pm$ 0.0034
ISRUC_S3 5-Class Dual 0.7671 $\pm$ 0.0091 0.8202 $\pm$ 0.0071 0.7856 $\pm$ 0.0031
Temporal 0.7314 $\pm$ 0.0210 0.7916 $\pm$ 0.0181 0.7565 $\pm$ 0.0244
Frequency 0.7390 $\pm$ 0.0601 0.7986 $\pm$ 0.0514 0.7728 $\pm$ 0.0361
Mixed 0.7400 $\pm$ 0.0217 0.7999 $\pm$ 0.0171 0.7673 $\pm$ 0.0157
Dual 0.7558 $\pm$ 0.0333 0.8130 $\pm$ 0.0264 0.7801 $\pm$ 0.0132
Dual 0.7359 $\pm$ 0.0324 0.7950 $\pm$ 0.0259 0.7621 $\pm$ 0.0311
Figure 3: Model and training data scaling laws of CodeBrain across three datasets on Cohen’s Kappa.
4.5 Vector Visualization
Figure 4: Decoupled time-frequency codes visualization on ISRUC_S3 dataset.
To better understand how the proposed TFDual-Tokenizer overcomes the limitations of single-codebook quantization in handling heterogeneous EEG signals, we visualize its learned temporal and frequency codes on the ISRUC_S3 dataset. As shown in Figure 4 (a) and (b), both temporal and frequency codebooks exhibit clustering behavior, capturing domain-specific features, such as characteristic waveforms or frequency patterns associated with specific sleep stages (e.g., N3). Figure 4 (c) and (d) further demonstrate that our decoupled design enables flexible composition: the same temporal code can pair with different frequency codes to represent different sleep stages, and vice versa. This indicates that temporal and frequency representations contribute both independently and complementarily to the final task representation. Such tokenizer supports a richer and more structured representation space for heterogeneous features, improving both performance and interpretability.
5 Conclusion
In this work, we propose CodeBrain, an efficient EEG foundation model that integrates three key components: a temporal–frequency decoupled tokenizer for heterogeneous representation learning, a structured state space backbone for modeling sparse global dependencies, and a sliding window attention mechanism for capturing local intra-patch patterns. Through comprehensive evaluation, CodeBrain consistently outperforms prior foundation models under linear probing, demonstrating strong generalization and transferability. Our results highlight the potential of CodeBrain as a versatile backbone for EEG-based brain–computer interface applications, offering a scalable and interpretable foundation for future developments in neural decoding and AI in neuroscience.
Appendix A Preliminaries
Convolution State Space Models
The state-space model is a classic model in control theory, and it represents the operational state of a system using first-order differential equations (ODE). A continuous state-space model can be defined in the following form:
$x^{\prime}(t)=Ax(t)+Bu(t),y(t)=Cx^{\prime}(t)+Du(t),$ (17)
where $u(t)$ is a vector that represents the input of the system, while $y(t)$ is a vector that represents the output of the system. $x(t)$ and its derivative $x^{\prime}(t)$ represent the latent states of the system, typically in the form of an N-D vector. And $A,B,C,D$ here are the state, input, output, and feedforward matrices, defining the relationship between the input, output, and state vector. Following Gu et al. [75], $\bar{D}$ is set equal to 0 since it can be replaced by the residual connection. Now, Equation 18 resembles an architecture similar to RNN, allowing us to recurrently compute $x_{k}$. Let the initial state be $x_{k-1}=0$, and we can unroll Equation 18 as follows:
$y_{k}=\overline{CA^{k}B}u_{0}+\overline{CA^{k-1}B}u_{1}+...+\overline{CAB}u_{k-1}+\overline{CB}u_{k}$ (18)
$y=\overline{K}u,\quad\overline{K}=(\overline{CB},\overline{CA^{1}B},...,\overline{CA^{k}B}).$ (19)
Therefore, SSM can be transformed from the form of a recurrent neural network to a convolutional neural network. During training, the $\overline{K}$ can be considered as a 1-D globe convolution kernel, so the $y$ can be calculated via the "long" convolution, allowing us to use the fast Fourier transform to efficiently compute the SSM convolutional kernel $\overline{K}$. However, directly computing the convolution in Equation 19 can be very expensive for long sequences. We can use the Fast Fourier Transform to accelerate it. The form of this convolution can be written as:
$y=F^{-1}_{N}D_{k}F_{N}u,D_{k}=\text{diag}(\overline{K}F_{N}),$ (20)
where $F_{N}$ denotes the DFT matrix of size $N$. This FFT convolution has a computational complexity of $O(nlog(n))$. Following [38, 76], we hope to parameterize $K$ directly rather than through $\{A,B,C\}$ because we can eliminate complex parameterization and accelerate the entire convolution.
Appendix B Pretraining Results
Our model follows a two-stage pretraining framework. In the first stage, we train the TFDual-Tokenizer, which independently tokenizes EEG signals in both the temporal and frequency domains. This tokenizer is optimized to reconstruct the original raw EEG signals, amplitude, and phase components, thereby producing discrete code representations with structural interpretability. In the second stage, we pretrain the EEGSSM encoder using a masked modeling objective: given the original EEG signals as input, the model learns to predict the corresponding masked tokens generated from the TFDual-Tokenizer.
This section reports the pretraining results of both stages, including loss convergence, reconstruction dynamics, and codebook utilization patterns.
B.1 TFDual-Tokenizer Pretraining Results
Total Training Loss
The total pretraining loss curve of the TFDual-Tokenizer is shown in Figure 5. The model demonstrates a rapid initial decrease in loss during the first few epochs, followed by a slower but consistent decline.
Figure 5: Pretraining Loss Curve of TFDual-Tokenizer.
Reconstruction Loss.
We report the pretraining reconstruction loss of the TFDual-Tokenizer in Figure 6. The temporal codebook is trained to reconstruct raw EEG signals in the time domain, while the frequency codebook is trained to reconstruct the corresponding amplitude and phase components in the frequency domain. All three loss curves exhibit a sharp initial decline, followed by a gradual convergence, indicating stable optimization.
Figure 6: Pretraining Loss Curve of TFDual-Tokenizer.
Unused Codes Analysis.
During pretraining, we track the number of unused codes in both the temporal and frequency codebooks of the TFDual-Tokenizer, each with a size of 4096. As shown in Figure 7, the frequency codebook demonstrates a rapid decrease in unused codes, while the temporal codebook shows a slower and more incremental reduction. A more detailed analysis of the temporal–frequency complementarity is provided in Section C.
Figure 7: Unused code dynamics of the TFDual-Tokenizer.
B.2 EEGSSSM Pretraining Results
We plot the pretraining loss curve of EEGSSM in Figure 8. We select epoch 10 as the checkpoint for downstream fine-tuning. We observe that the pretraining loss of EEGSSM decreases rapidly from epoch 1 to 6 (9.04 → 6.39), then flattens gradually after epoch 10 (6.01 → 5.66). Using epoch 10 for fine-tuning is a balance between representation strength and generalization. Overtraining on EEG, prone to noise and inter-subject variability, can reduce transferability. Epoch 10 serves as a conservative yet effective checkpoint. This practice is consistent with trends in foundation models from NLP [17] and vision [77], where mid-training checkpoints often lead to better downstream performance than final ones due to reduced overfitting to the pretext task.
Figure 8: Pretraining Loss Curve of EEGSSM.
Appendix C Class-Specific Token Ratio Analysis for TFDual-Tokenizer
To better understand the contribution of the TFDual-Tokenizer module, which decouples time and frequency domain representations, in learning discriminative EEG features, we conduct an analysis of class-specific tokens. A token (code) is considered class-specific if it predominantly appears in samples of a single class. Formally, for a given token $c$, let $N_{c}^{(y)}$ denote the number of times $c$ appears in class $y$. The dominance ratio is defined as:
$\text{Dominance}(c)=\frac{\max_{y}N_{c}^{(y)}}{\sum_{y}N_{c}^{(y)}}$ (21)
If $\text{Dominance}(c)>\tau$ (we use $\tau=1$), the token is deemed class-specific. The class-specific ratio for a codebook is then computed as:
[…]
Figure 9 illustrates the proportion of class-specific tokens derived from three configurations: using only the temporal codebook, only the frequency codebook, and a combination of both (TF-Decoupled). We employ two independent codebooks to capture complementary information via the proposed TFDual-Tokenizer module.
Across all four datasets, including two for emotion recognition (FACED and SEED-V) and two for sleep staging (ISRUC_S3 and ISRUC_S1), the decoupled codebook consistently achieves the highest class-specific token ratios, reaching 54.7% on ISRUC_S3 and 46.4% on FACED. These results confirm that decoupling temporal and frequency domain information significantly enhances the model’s ability to capture class-discriminative patterns.
Temporal and frequency domains exhibit inherently heterogeneous characteristics: temporal signals capture transient waveform dynamics, while frequency components reflect oscillatory structures over broader time scales. By decoupling these two domains and learning separate token representations, the model is able to preserve their distinct informational content. The TFDual-Tokenizer then integrates temporally localized patterns—such as emotion-related bursts or sleep-stage-specific oscillations—that are highly class-discriminative. The TFDual-Tokenizer enables richer and more effective EEG representations.
Figure 9: Class-Specific Code Ratio Across Different Codebooks.
Appendix D Improving Temporal Codebook Learning via Contrastive Loss
To improve the learning of the temporal codebook in our TFDual-Tokenizer, we introduce a contrastive loss as one of the objectives during pretraining. This design is motivated by observations from prior work LaBraM [20], where the authors report that reconstructing raw EEG signals leads to unconvergence, and thus omit the temporal reconstruction objective entirely.
To better understand this limitation, we first implemented a baseline reconstruction of raw signals within the LaBraM framework and observed that the training loss plateaued at a high value (between 0.128 and 0.131), showing no convergence over training epochs. To mitigate this, we introduce a TFConv module before the Transformer encoder, designed to extract temporal-frequency representations before tokenization. While this stabilizes training to some extent, we still observe significant issues with code utilization and loss convergence. Therefore, we incorporate a lightweight contrastive loss, applied over temporal representations before quantization, to encourage the model to organize similar input patterns closer in the latent space. As shown in Figure 10, this improves the optimization of the reconstruction loss and reduces the number of unused temporal codes during training.
These results demonstrate that contrastive regularization acts as an effective prior for stabilizing discrete token learning, particularly when reconstructing raw signals. It both improves convergence and mitigates codebook collapse in the temporal branch of the tokenizer.
Figure 10: Effect of Contrastive Loss on Temporal Codebook Learning.
Appendix E Dataset Description
We evaluate the CodeBrain across eight diverse downstream tasks covering ten publicly available EEG datasets. Notably, these datasets exhibit substantial variability in channel configurations (ranging from 6 to 64), sequence lengths (from 1s to 30s), and task complexities (2-class to 9-class), highlighting the versatility and robustness of CodeBrain across different EEG applications. The following sections describe each dataset in detail, including its task objective and data split strategy. A comprehensive analysis of each dataset is provided below.
Emotion Recognition.
We conduct emotion recognition experiments on two widely used EEG datasets: FACED [68] and SEED-V [69].
The FACED dataset (Finer-Grained Affective Computing EEG Dataset) is a large-scale EEG dataset proposed by Chen et al. [68] for emotion recognition tasks. It consists of 32-channel EEG recordings sampled at 250 Hz from 123 participants, each exposed to 28 video clips designed to elicit nine distinct emotional states: amusement, inspiration, joy, tenderness, anger, fear, disgust, sadness, and neutral. These cover both positive and negative affective categories. Each EEG trial is 10 seconds long and is subsequently resampled to 200 Hz, resulting in a total of 10,332 clean EEG segments. For fair comparison, we adopt the same subject-wise split as in [19]: subjects 1–80 are used for training, 81–100 for validation, and 101–123 for testing, ensuring no subject overlap across splits and enabling evaluation of cross-subject generalization.
SEED-V [69] is an EEG dataset designed for emotion recognition, covering five emotional categories: happy, sad, neutral, disgust, and fear. It consists of 62-channel EEG recordings collected at 1000 Hz from 16 subjects, each participating in three sessions. Each session includes 15 trials, which are evenly divided into training, validation, and test sets (5 trials each). The EEG signals are segmented into 1-second windows, yielding a total of 117,744 samples, and resampled to 200 Hz for consistency. The dataset provides rich temporal structure and inter-subject variability, making it a strong benchmark for evaluating generalization in emotion-related EEG modeling.
Sleep Staging.
We use two datasets, ISRUC_S1 and ISRUC_S3 [70], for sleep stage classification. Both datasets are annotated according to the American Academy of Sleep Medicine (AASM) standard [78], with five sleep stages: Wake, NREM1 (N1), NREM (N2), NREM (N3), and REM. Each EEG segment corresponds to a 30-second epoch.
ISRUC_S1 includes EEG recordings from 100 subjects using six channels at a sampling rate of 200 Hz. We adopt a subject-wise split, with 80 subjects for training, 10 for validation, and 10 for testing. As the transition rules between sleep stages carry important temporal patterns, we follow prior work [18, 19] and insert a Transformer layer on top of the projection head during fine-tuning to better capture sequence-level dependencies. We set the input sequence length to 20, and discard segments that cannot be evenly divided. In total, 86,320 labeled samples are retained.
ISRUC_S3 is a smaller dataset comprising recordings from 10 subjects, also sampled at 200 Hz with six channels, totaling 8,500 labeled segments. We follow an 8:1:1 subject-wise split for training, validation, and testing.
Imagined Speech Classification.
The BCIC2020-T3 dataset [71] was released as part of the 2020 International Brain–Computer Interface Competition and focuses on imagined speech decoding. It contains EEG recordings from 15 participants who were instructed to silently imagine speaking five specific words or phrases, "hello", "help me", "stop", "thank you", and "yes". EEG signals were collected using 64 scalp channels at a sampling rate of 256 Hz and were subsequently resampled to 200 Hz for preprocessing consistency. Each subject completed 80 trials per class, resulting in a total of 6,000 trials. The dataset provides predefined training, validation, and test splits, with 60, 10, and 10 trials per class, respectively, facilitating fair model evaluation like existing baselines [19].
Mental Stress Detection.
The Mental Arithmetic dataset [73] supports the task of mental stress detection using EEG signals. It contains recordings from 36 subjects under two distinct cognitive conditions: resting and active engagement in mental arithmetic. EEG data labeled as “no stress” correspond to resting periods prior to the task, while “stress” labels are assigned to recordings during task performance. The signals were acquired using 20 electrodes placed according to the international 10–20 system, with an original sampling rate of 500 Hz. For consistency, the signals are resampled to 200 Hz and band-pass filtered between 0.5–45 Hz to suppress noise. Each recording is segmented into 5-second windows, yielding a total of 1,707 samples. We adopt a subject-wise split for fair evaluation with existing baselines [19]: subjects 1–28 for training, 29–32 for validation, and 33–36 for testing.
Seizure Detection.
The CHB-MIT dataset [72] is a widely used benchmark for seizure detection from EEG signals. It contains long-term EEG recordings from 23 patients diagnosed with intractable epilepsy, collected at the Children’s Hospital Boston. The subjects underwent continuous monitoring over several days, during which seizures were recorded following the tapering of anti-epileptic medications. EEG signals were acquired using the international 10–20 system and originally sampled at 256 Hz. In our setting, we adopt 16 channels commonly used in prior work [29, 19], resample all signals to 200 Hz, and segment them into 10-second non-overlapping windows, yielding 326,993 labeled samples across seizure and non-seizure classes. We follow a subject-wise split: subjects 1–19 for training, 20–21 for validation, and 22–23 for testing. Notably, this dataset is highly imbalanced, with seizure events constituting only a small fraction of the total samples, posing significant challenges for model training and evaluation.
Motor Imagery Classification.
The SHU-MI dataset [74] is designed for binary motor imagery classification, where participants are instructed to imagine movements of either the left or right hand. EEG signals were recorded from 25 subjects using a 32-channel setup at an original sampling rate of 250 Hz. To ensure consistency with the pre-training setting, all signals are resampled to 200 Hz and segmented into 4-second non-overlapping windows, resulting in 11,988 labeled samples. A subject-wise split is applied for fair model evaluation like existing baselines [19], with subjects 1–15 used for training, 16–20 for validation, and 21–25 for testing. This dataset supports the development of BCI systems that decode motor intentions from brain activity without actual movement.
Event Type Classification.
The TUEV dataset [66] is a clinically annotated EEG corpus used for multi-class event type classification. It includes six event categories: spike and sharp wave (SPSW), generalized periodic epileptiform discharges (GPED), periodic lateralized epileptiform discharges (PLED), eye movements (EYEM), artifacts (ARTF), and background activity (BCKG). EEG signals were originally recorded at 256 Hz using 23 channels. In line with prior work [19], we preprocess the data by selecting 16 bipolar montage channels based on the international 10–20 system. The signals are band-pass filtered between 0.3–75 Hz to suppress low- and high-frequency noise, and a 60 Hz notch filter is applied to eliminate power line interference. All recordings are resampled to 200 Hz and segmented into 5-second windows, yielding 112,491 labeled samples. We follow the official train-test split and further divide the training subjects into training and validation sets in an 8:2 ratio, consistent with established benchmarks.
Abnormal Detection.
The TUAB dataset [66] is employed for binary abnormal EEG detection, where each EEG recording is labeled as either normal or abnormal based on clinical interpretation. Originally recorded at 256 Hz using 23 channels, the dataset provides large-scale EEG recordings suitable for evaluating diagnostic models. To ensure fair comparison with prior work [19], we follow a similar preprocessing protocol. Specifically, we select 16 bipolar montage channels following the international 10–20 system, apply band-pass filtering between 0.3–75 Hz to eliminate low- and high-frequency artifacts, and remove 60 Hz power line interference using a notch filter. The EEG signals are then resampled to 200 Hz and segmented into 10-second windows, resulting in 409,455 labeled samples. We follow the official train-test split and further divide the training set into training and validation subsets using an 8:2 subject-wise ratio, consistent with existing benchmarks.
Appendix F Baselines and Metrics Description
F.1 Metrics
To comprehensively evaluate our model, we compare it with a set of strong baselines commonly used in EEG analysis. These baselines are evaluated using metrics tailored for class-imbalanced scenarios, which are prevalent in EEG datasets. The metrics include:
  • Balanced Accuracy, which averages the recall across all classes and is particularly suitable for imbalanced multi-class classification tasks.
  • AUROC and AUC-PR, which assess the performance of binary classifiers under different thresholds. While AUROC measures the trade-off between sensitivity and specificity, AUC-PR focuses on precision-recall trade-offs, especially informative under severe class imbalance.
  • Cohen’s Kappa, which quantifies inter-class agreement beyond chance and is employed as the primary metric for multi-class classification.
  • Weighted F1 Score, which combines precision and recall while adjusting for class support, ensuring fair performance measurement across imbalanced datasets.
For model selection and comparison, AUROC is used as the main evaluation metric for binary classification tasks, and Cohen’s Kappa is used for multi-class scenarios.
F.2 Baselines
We compare our CodeBrain model against 5 publicly available EFMs that have released pre-trained weights, covering a diverse set of pretraining strategies.
We do not include traditional supervised baselines such as EEGNet [79] or ST-Transformer [80] in this comparison, as previous studies have consistently shown that EFMs significantly outperform these smaller supervised models. Our focus is therefore on evaluating the effectiveness of different foundation model designs and pretraining paradigms under comparable settings.
BENDR [47]: We adopted BENDR (Bert-inspired Neural Data Representations) as our baseline model, as introduced by Kostas et al. BENDR is a pioneering deep learning architecture for Electroencephalography (EEG) data, leveraging transformers and a contrastive self-supervised learning task. This approach enables the model to learn meaningful representations from vast amounts of unlabeled EEG data.
BIOT [29]: BIOT (Biosignal Transformer for Cross-data Learning in the Wild) is a transformer-based architecture designed to handle cross-dataset EEG signal classification under domain shifts. It leverages a domain-invariant attention mechanism and contrastive representation learning to enhance generalization across different recording conditions and subject populations.
LaBraM[20]: LaBraM (Large Brain Model) proposes a scalable transformer-based framework designed to learn generic EEG representations from large-scale brain signal datasets. By pretraining on a diverse corpus of EEG recordings, the model captures rich temporal and spatial features that transfer effectively to various downstream BCI tasks. The architecture incorporates efficient self-attention mechanisms and task-specific adapters to support flexible fine-tuning.
EEGPT [18]: EEGPT employs a dual self-supervised learning strategy that combines masked autoencoding with spatio-temporal representation alignment, enhancing feature quality by focusing on high signal-to-noise ratio (SNR) representations rather than raw signals. The model’s hierarchical architecture decouples spatial and temporal processing, improving computational efficiency and adaptability to various brain-computer interface (BCI) applications.
CBraMod [19]: CBraMod (Criss-Cross Brain Foundation Model) is a transformer-based EEG foundation model that addresses the heterogeneous spatial and temporal dependencies inherent in EEG signals. It introduces a criss-cross transformer architecture comprising parallel spatial and temporal attention mechanisms, enabling separate yet simultaneous modeling of spatial and temporal relationships.
Appendix G Hyperparameter Setting
We provide detailed hyperparameter configurations for the two-stage pretraining of our CodeBrain model and the fine-tuning settings across ten downstream tasks.
G.1 Pretraining Settings
The pretraining process consists of two stages:
  1. 1.
    training the TFDual-Tokenizer
  2. 2.
    training the EEGSSM framework
The hyperparameters used in each stage are summarized in Table 4 and Table 5, respectively.
Table 4: Hyperparameters for TFDual-Tokenizer.
Hyperparameters Values
TFConv
   Input channels {1, 8, 4}
   Output channels {8, 4, 4}
   Kernel size {(1, 15), (1, 3), (1, 3)}
   Stride {(1, 8), (1, 1), (1, 1)}
   Padding {(0, 7), (0, 1), (0, 1)}
Transformer encoder layers 12
Transformer decoder layers 3
Hidden size 200
MLP size 800
Attention head number 8
Temporal Codebook size 4096 $\times$ 32
Frequency Codebook size 4096 $\times$ 32
Batch size 256
Peak learning rate 1e-4
Minimal learning rate 1e-5
Learning rate scheduler Cosine
Optimizer AdamW
Adam $\beta$ (0.9, 0.99)
Weight decay 1e-4
Total epochs 20
Data stride 200
Table 5: Hyperparameters for EEGSSM.
Hyperparameters Values
Epochs 10
Batch size 256
Dropout 0.1
Optimizer Adam
Learning rate 1e-4
Adam $\beta$ (0.9, 0.999)
Adam $\epsilon$ 1e-8
Weight decay 5e-3
Scheduler CosineAnnealingLR
Minimal learning rate 1e-5
Clipping gradient norm 5
G.2 Fine-tuning Settings on Downstream Tasks
The CodeBrain model is fine-tuned on ten downstream EEG classification tasks, each with task-specific hyperparameters. Following the general strategy adopted by prior EFMs, we adopt a lightweight three-layer MLP as the probe head for all downstream tasks and fine-tune the entire model end-to-end. Table 6 lists the fine-tuning configurations including learning rate, weight decay, dropout rate, and batch size for each task. For sleep staging tasks, due to their strong temporal structure, we follow prior work [18, 19] and insert an additional Transformer encoder on top of the projection head to jointly model the sequence of 20 consecutive EEG segments. This enables the model to capture inter-epoch transitions critical to sleep stage classification.
Table 6: Fine-tuning Hyperparameters for Downstream Tasks.
Dataset Learning Rate Weight Decay Dropout Batch Size
FACED 5e-5 5e-4 0.1 16
SEED-V 5e-5 1e-2 0.1 64
ISRUC_S1 1e-4 1e-1 0.2 48
ISRUC_S3 1e-4 1e-1 0.2 48
BCIC2020-T3 5e-5 5e-2 0.1 32
Mental Arithmetic 3e-5 1e-3 0.1 32
CHB-MIT 3e-5 1e-2 0.4 64
SHU-MI 5e-5 5e-3 0.3 64
TUEV 2e-5 5e-4 0.3 64
TUAB 1e-5 5e-5 0.4 512
Appendix H Model Comparison
Backbone Efficiency Comparison
To evaluate the computational efficiency of our proposed SGConv module, we conduct an ablation study by replacing it with three common sequence modeling modules: CNN, LSTM, and Transformer. We compare their model sizes, floating-point operations (FLOPs), and iteration times, as shown in Figure 11. Specifically, the CNN variant uses a three-layer depthwise separable convolution block, while the LSTM and Transformer variants use a single layer of standard LSTM and Transformer Encoder (implemented by Pytorch), respectively. In terms of parameter count, SGConv contains only 8.74M parameters, significantly fewer than CNN (28.22M), LSTM (30.35M), and Transformer (27.79M). For FLOPs, SGConv also achieves the lowest computational cost at 15.17G, compared to Transformer’s 21.2G. Regarding iteration time, SGConv is slightly slower than Transformer and CNN models in terms of training speed, but it outperforms the LSTM model. In summary, SGConv effectively reduces the number of parameters while maintaining computational complexity, which helps the model to be trained and inferred on smaller GPUs.
Figure 11: Computational Overhead of Using Different Backbones in the EEGSSM Module.
Additional Evaluation on Other BCI Tasks
We report the performance of CodeBrain on four additional EEG datasets not included in the main text, covering diverse domains of sleep staging, motor imagery, event detection, and abnormality classification in Tables 7 to 10. These allow us to assess the cross-domain generalization ability of our pretrained model beyond the main text.
We note that both TUAB and TUEV originate from the TUH EEG corpus [66], which overlaps with our pretraining source (TUEG). To avoid overfitting to this distribution and promote generalization, we stop pretraining at epoch 10 as discussed in Section B.2. While this may limit gains on TUH datasets compared to previous EFM, such as CBraMod (trained for 40 epochs in the same pretraining dataset) [19], CodeBrain still achieves superior or competitive results.
ISRUC_S1
As shown in Table 7, CodeBrain achieves state-of-the-art performance on ISRUC_S1 in terms of Cohen’s Kappa (0.7476) and Weighted F1 (0.8020), slightly surpassing CBraMod[19] by +0.34 and +0.09 points, respectively. Its Balanced Accuracy of 0.7835 is also competitive, trailing the best result by only -0.30. These results highlight the model’s ability to capture temporal dependencies and learn discriminative representations for 5-class sleep staging under a cross-subject setting.
Table 7: Performance Comparison on the ISURC_S1 (5-Class) dataset.
Methods Cohen’s Kappa Weighted F1 Balanced Accuracy
BENDR[47] 0.6956 ± 0.0053 0.7569 ± 0.0049 0.7401 ± 0.0056
BIOT[29] 0.7192 ± 0.0231 0.7790 ± 0.0146 0.7527 ± 0.0121
LaBraM[20] 0.7231 ± 0.0182 0.7810 ± 0.0133 0.7633 ± 0.0102
EEGPT[18] 0.2223 ± 0.0227 0.3111 ± 0.0110 0.4012 ± 0.0177
CBraMod[19] 0.7442 ± 0.0152 0.8011 ± 0.0099 0.7865 ± 0.0110
CodeBrain 0.7476 ± 0.0040 0.8020 ± 0.0018 0.7835 ± 0.0033
SHU-MI
As shown in Table 8, CodeBrain achieves the best overall performance on SHU-MI across all three metrics. It obtains an AUROC of 0.7124 and an AUC-PR of 0.7166, slightly improving over the previous best by +1.36 and +0.27 points, respectively. For Balanced Accuracy, it reaches 0.6431 (+0.61), with notably lower variance. These results underscore its strong generalization to motor imagery decoding under a cross-subject protocol.
Table 8: Performance Comparison on the SHU-MI (2-Class) dataset.
Methods AUROC AUC-PR Balanced Accuracy
BENDR[47] 0.5863 ± 0.0280 0.5853 ± 0.0268 0.5573 ± 0.0227
BIOT[29] 0.6609 ± 0.0127 0.6770 ± 0.0119 0.6179 ± 0.0183
LaBraM[20] 0.6604 ± 0.0091 0.6761 ± 0.0083 0.6166 ± 0.0192
EEGPT[18] 0.6241 ± 0.0071 0.6266 ± 0.0133 0.5778 ± 0.0162
CBraMod[19] 0.6988 ± 0.0068 0.7139 ± 0.0088 0.6370 ± 0.0151
CodeBrain 0.7124 ± 0.0050 0.7166 ± 0.0106 0.6431 ± 0.0066
TUEV
As shown in Table 9, CodeBrain achieves the highest Cohen’s Kappa (0.6912, an improvement of +0.0140 over the best baseline [19]) and Weighted F1 (0.8362) on TUEV. Although its Balanced Accuracy is lower than CBraMod [19], we attribute this to reduced sensitivity on the rare SPSW class. Since TUEV shares distributional overlap with our pretraining source (TUEG), we stop pretraining at epoch 10 to prevent overfitting, unlike CBraMod’s 40-epoch training as discussed in subsection B.2. We also report LaBraM’s results based on its original 23-channel setting [20], while CodeBrain follows the 16-channel configuration used in CBraMod. Similarly, EEGPT [18] does not adopt a linear fine-tuning protocol but applies two convolutional layers before entering the foundation model, followed by an MLP head. While such architectural choices may enhance performance, we follow their respective fine-tuning settings. In such cases, CodeBrain still outperforming both LaBraM and EEGPT under their own fine-tuning settings clearly demonstrates the robustness of our approach.
Table 9: Performance Comparison on the TUEV (6-Class) dataset.
Methods Cohen’s Kappa Weighted F1 Balanced Accuracy
BENDR[47] 0.4271 ± 0.0238 0.6755 ± 0.0216 0.4363 ± 0.0245
BIOT[29] 0.5273 ± 0.0249 0.7492 ± 0.0082 0.5281 ± 0.0225
LaBraM[20] 0.6637 ± 0.0093 0.8312 ± 0.0052 0.6409 ± 0.0065
EEGPT[18] 0.6351 ± 0.0134 0.8187 ± 0.0063 0.6232 ± 0.0114
CBraMod[19] 0.6772 ± 0.0096 0.8342 ± 0.0064 0.6671 ± 0.0107
CodeBrain 0.6912 ± 0.0101 0.8362 ± 0.0048 0.6428 ± 0.0062
TUAB
As shown in Table 10, CodeBrain achieves the highest Balanced Accuracy (0.8294) on TUAB, slightly outperforming CBraMod [19]. Similar to TUEV, TUAB is part of the TUH EEG corpus family [66] and thus closely aligned with our pretraining source (TUEG). As discussed in subsection B.2, we adopt an early stopping strategy at epoch 10 to mitigate overfitting to this distribution, which may partly account for the slightly lower AUROC and AUC-PR compared to CBraMod, trained for 40 epochs on the same dataset. While LaBraM leverages a 23-channel montage [20] and EEGPT [18] employs two convolutional layers before the foundation model, we retain their respective fine-tuning protocols for comparison. Despite these potentially stronger configurations, CodeBrain still exceeds both models under their own settings, highlighting its strong and consistent generalization.
Table 10: Performance Comparison on the TUAB (2-Class) dataset.
Methods Balanced Accuracy AUC-PR AUROC
BENDR[47] 0.7714 ± 0.0248 0.8412 ± 0.0215 0.8426 ± 0.0237
BIOT[29] 0.7959 ± 0.0057 0.8792 ± 0.0023 0.8815 ± 0.0043
LaBraM[20] 0.8140 ± 0.0019 0.8965 ± 0.0016 0.9022 ± 0.0009
EEGPT[18] 0.8038 ± 0.0040 0.8891 ± 0.0018 0.8811 ± 0.0015
CBraMod[19] 0.8289 ± 0.0022 0.9258 ± 0.0008 0.9227 ± 0.0011
CodeBrain 0.8294 ± 0.0013 0.9100 ± 0.0006 0.9030 ± 0.0009
Appendix I Ablation on Mask Ratio
We conduct an ablation study to investigate the effect of the mask ratio in the EEGSSM pretraining framework. As shown in Tables 1113 and Figure 12, downstream performance consistently exhibits a U-shaped trend with respect to the masking ratio across all three datasets: FACED, SEED-V, and ISRUC_S3. Moderate masking (e.g., ratios around 0.4–0.6) leads to optimal performance, whereas excessively low (e.g., 0.1) or high (e.g., 0.9) ratios degrade generalization.
Table 11: Performance of CodeBrain on FACED Dataset under Different Mask Ratios.
Mask Ratio Cohen’s Kappa Weighted F1 Balanced Accuracy
0.1 0.5056 ± 0.0054 0.5652 ± 0.0063 0.5579 ± 0.0059
0.2 0.5111 ± 0.0042 0.5721 ± 0.0030 0.5692 ± 0.0035
0.3 0.5196 ± 0.0081 0.5757 ± 0.0059 0.5707 ± 0.0072
0.4 0.5253 ± 0.0042 0.5831 ± 0.0039 0.5795 ± 0.0046
0.5 0.5270 ± 0.0039 0.5849 ± 0.0033 0.5816 ± 0.0028
0.6 0.5156 ± 0.0066 0.5720 ± 0.0075 0.5671 ± 0.0095
0.7 0.5113 ± 0.0087 0.5699 ± 0.0063 0.5621 ± 0.0075
0.8 0.5026 ± 0.0057 0.5467 ± 0.0044 0.5408 ± 0.0042
0.9 0.4906 ± 0.0083 0.5369 ± 0.0084 0.5345 ± 0.0095
Table 12: Performance of CodeBrain on SEED-V Dataset under Different Mask Ratios.
Mask Ratio Cohen’s Kappa Weighted F1 Balanced Accuracy
0.1 0.2523 ± 0.0051 0.4081 ± 0.0030 0.3968 ± 0.0031
0.2 0.2633 ± 0.0033 0.4137 ± 0.0027 0.4071 ± 0.0035
0.3 0.2703 ± 0.0043 0.4200 ± 0.0048 0.4121 ± 0.0041
0.4 0.2734 ± 0.0042 0.4244 ± 0.0029 0.4142 ± 0.0055
0.5 0.2735 ± 0.0032 0.4235 ± 0.0022 0.4137 ± 0.0023
0.6 0.2699 ± 0.0041 0.4207 ± 0.0059 0.4106 ± 0.0048
0.7 0.2642 ± 0.0045 0.4158 ± 0.0033 0.4091 ± 0.0040
0.8 0.2603 ± 0.0031 0.4116 ± 0.0041 0.4009 ± 0.0025
0.9 0.2534 ± 0.0036 0.4085 ± 0.0024 0.3993 ± 0.0055
Table 13: Performance of CodeBrain on ISRUC_S3 Dataset under Different Mask Ratios.
Mask Ratio Cohen’s Kappa Weighted F1 Balanced Accuracy
0.1 0.7252 ± 0.0069 0.7865 ± 0.0039 0.7686 ± 0.0072
0.2 0.7403 ± 0.0023 0.7982 ± 0.0026 0.7752 ± 0.0032
0.3 0.7501 ± 0.0057 0.8013 ± 0.0062 0.7766 ± 0.0020
0.4 0.7608 ± 0.0060 0.8168 ± 0.0078 0.7846 ± 0.0062
0.5 0.7671 ± 0.0091 0.8202 ± 0.0071 0.7856 ± 0.0031
0.6 0.7661 ± 0.0068 0.8219 ± 0.0052 0.7826 ± 0.0076
0.7 0.7577 ± 0.0068 0.8161 ± 0.0040 0.7782 ± 0.0061
0.8 0.7471 ± 0.0031 0.8090 ± 0.0041 0.7658 ± 0.0025
0.9 0.7504 ± 0.0078 0.8125 ± 0.0063 0.7668 ± 0.0065
Figure 12: Performance Across Different Mask Ratios on FACED, SEED-V, and ISRUC_S3.
Figure 13: EEGSSM Pre-Training Loss Curve for Different Mask Ratios.
To further illustrate this pattern, we visualize the training loss curves across different mask ratios in Figure 13. Interestingly, higher mask ratios result in slower convergence and higher final training loss, which is expected due to the increased difficulty of the reconstruction task. In contrast, lower mask ratios lead to faster and smoother loss reduction, but do not necessarily yield better downstream performance. This observation suggests a possible optimization-vs-generalization trade-off: easier pretext tasks (low mask ratio) are more optimizable but may encourage the model to learn shortcut solutions with limited generalizability, while overly difficult tasks (high mask ratio) may hinder effective representation learning due to insufficient learning signal. Moderate masking strikes a balance by being sufficiently challenging to promote abstraction, while still being learnable, thereby facilitating better generalization across downstream tasks.
Appendix J Detailed Results on Scaling Laws
We provide the detailed scaling law results for both data and model size across three representative EEG datasets (FACED, SEED-V, and ISRUC_S3), covering three evaluation metrics. For brevity, only Cohen’s kappa scores are included in the main text, while full results are provided in this section. Prior work [19, 20] has explored the effect of scaling EEG foundation models using 1 to 1000 hours of pretraining data and model sizes up to 5M parameters. We extend this analysis in two key dimensions:
  1. 1.
    Scaling the pretraining data volume from 1k up to 9k hours.
  2. 2.
    Investigating model scaling by varying the depth of the EEGSSM encoder from 3 layers (6.82M parameters) to 8 layers (15.17M parameters).
J.1 Scaling Laws with Respect to Training Data Volume
We examine how the volume of pretraining data influences the downstream performance of CodeBrain. Specifically, we scale the pretraining duration from 1k to 9k hours and evaluate the resulting models on three downstream datasets: FACED, SEED-V, and ISRUC_S3. Detailed quantitative results across three evaluation metrics (Cohen’s kappa, weighted F1 score, and balanced accuracy) are presented in Tables 1415, and 16, respectively. As shown in Figure 14, increasing the amount of pretraining data generally leads to consistent improvements across all datasets and metrics. On FACED and ISRUC_S3, performance gains are steady throughout the entire range up to 9k hours, while on SEED-V, the trend is more modest and plateaus after 5k hours. These results highlight the importance of large-scale data for representation learning in EEG and suggest that further scaling may continue to yield performance benefits.
Table 14: Training Data Scaling Laws of CodeBrain on FACED Dataset.
Training Data Cohen’s Kappa Weighted F1 Balanced Accuracy
1000 Hours 0.4883 ± 0.0102 0.5379 ± 0.0159 0.5344 ± 0.0150
2000 Hours 0.5001 ± 0.0084 0.5459 ± 0.0107 0.5412 ± 0.0134
3000 Hours 0.5058 ± 0.0075 0.5606 ± 0.0092 0.5574 ± 0.0086
4000 Hours 0.5084 ± 0.0018 0.5632 ± 0.0040 0.5600 ± 0.0034
5000 Hours 0.5034 ± 0.0069 0.5581 ± 0.0114 0.5549 ± 0.0087
6000 Hours 0.5136 ± 0.0105 0.5682 ± 0.0057 0.5650 ± 0.0071
7000 Hours 0.5196 ± 0.0114 0.5716 ± 0.0095 0.5694 ± 0.0103
8000 Hours 0.5201 ± 0.0087 0.5751 ± 0.0065 0.5719 ± 0.0077
9000 Hours 0.5270 ± 0.0039 0.5849 ± 0.0033 0.5816 ± 0.0028
Table 15: Training Data Scaling Laws of CodeBrain on SEED-V Dataset.
Training Data Cohen’s Kappa Weighted F1 Balanced Accuracy
1000 Hours 0.2584 ± 0.0044 0.3946 ± 0.0063 0.3799 ± 0.0096
2000 Hours 0.2648 ± 0.0062 0.4042 ± 0.0091 0.3961 ± 0.0084
3000 Hours 0.2689 ± 0.0055 0.4117 ± 0.0075 0.4028 ± 0.0082
4000 Hours 0.2672 ± 0.0043 0.4121 ± 0.0054 0.4022 ± 0.0049
5000 Hours 0.2678 ± 0.0069 0.4120 ± 0.0084 0.4026 ± 0.0067
6000 Hours 0.2669 ± 0.0031 0.4113 ± 0.0067 0.4030 ± 0.0102
7000 Hours 0.2686 ± 0.0043 0.4129 ± 0.0052 0.4027 ± 0.0048
8000 Hours 0.2703 ± 0.0049 0.4165 ± 0.0034 0.4094 ± 0.0054
9000 Hours 0.2735 ± 0.0032 0.4235 ± 0.0022 0.4137 ± 0.0023
Table 16: Training Data Scaling Laws of CodeBrain on the ISRUC_S3 Dataset.
Training Data Cohen’s Kappa Weighted F1 Balanced Accuracy
1000 Hours 0.7340 ± 0.0187 0.7826 ± 0.0154 0.7505 ± 0.0192
2000 Hours 0.7347 ± 0.0051 0.7869 ± 0.0075 0.7524 ± 0.0032
3000 Hours 0.7540 ± 0.0106 0.8012 ± 0.0093 0.7694 ± 0.0089
4000 Hours 0.7590 ± 0.0042 0.8108 ± 0.0074 0.7753 ± 0.0065
5000 Hours 0.7610 ± 0.0058 0.8124 ± 0.0102 0.7794 ± 0.0086
6000 Hours 0.7681 ± 0.0125 0.8190 ± 0.0100 0.7856 ± 0.0089
7000 Hours 0.7648 ± 0.0079 0.8170 ± 0.0076 0.7845 ± 0.0076
8000 Hours 0.7668 ± 0.0121 0.8182 ± 0.0109 0.7851 ± 0.0116
9000 Hours 0.7671 ± 0.0091 0.8202 ± 0.0071 0.7856 ± 0.0031
Figure 14: Training Data Scaling Laws on FACED, SEED-V, and ISRUC_S3.
In addition, we visualize the pretraining optimization behavior across different data scales in Figure 15. As expected, larger pretraining data consistently lead to lower training loss, indicating more effective representation learning. Notably, the convergence curves become progressively smoother and more stable as training data volume increases, suggesting improved optimization stability in large-scale regimes. While smaller training data volumes (e.g., 1k–3k hours) show relatively high starting loss and slower convergence, larger training data volumes (6k–9k hours) reach lower final losses and exhibit diminishing returns, aligning with trends observed in downstream performance. These findings provide further empirical support for the scalability of EEG foundation models and reinforce the role of large data in enhancing both optimization and generalization.
Figure 15: EEGSSM Pre-Training Loss Curve for Different Training Data Volume.
J.2 Scaling Laws with Respect to Model Size
We further investigate how model parameters affect downstream performance by scaling the number of layers in the EEGSSM encoder from 3 to 8, resulting in parameter counts ranging from 6.82M to 15.17M. Detailed results across the FACED, SEED-V, and ISRUC_S3 datasets are provided in Tables 1718, and 19, respectively. To visualize the trend more clearly, Figure 16 presents the performance curves as model size increases. Across all three datasets and evaluation metrics, we observe a consistent performance gain as the model size increases. The improvements are particularly pronounced on the FACED and ISRUC_S3 datasets, where all three metrics show steady growth up to the largest model. In contrast, performance on SEED-V improves more modestly and begins to plateau beyond 13.5M parameters. These results suggest that increasing model capacity can enhance generalization ability, especially for datasets with richer structure or more complex temporal dynamics, while also indicating that optimal scaling may be task-dependent.
Table 17: Model Size Scaling Laws of CodeBrain on the FACED Dataset.
Model Size Cohen’s Kappa Weighted F1 Balanced Accuracy
6.82M 0.4818 ± 0.0165 0.5362 ± 0.0113 0.5317 ± 0.0182
8.49M 0.4988 ± 0.0082 0.5506 ± 0.0091 0.5497 ± 0.0084
10.16M 0.5096 ± 0.0049 0.5686 ± 0.0104 0.5642 ± 0.0067
11.83M 0.5144 ± 0.0113 0.5705 ± 0.0085 0.5678 ± 0.0095
13.50M 0.5178 ± 0.0069 0.5772 ± 0.0080 0.5746 ± 0.0067
15.17M 0.5270 ± 0.0039 0.5849 ± 0.0033 0.5816 ± 0.0028
Table 18: Model Size Scaling Laws of CodeBrain on the SEED-V Dataset.
Model Size Cohen’s Kappa Weighted F1 Balanced Accuracy
6.82M 0.2609 ± 0.0078 0.4004 ± 0.0112 0.3956 ± 0.0098
8.49M 0.2638 ± 0.0080 0.4108 ± 0.0091 0.4030 ± 0.0121
10.16M 0.2645 ± 0.0102 0.4127 ± 0.0158 0.4013 ± 0.0084
11.83M 0.2663 ± 0.0056 0.4202 ± 0.0068 0.4079 ± 0.0055
13.50M 0.2724 ± 0.0057 0.4211 ± 0.0060 0.4120 ± 0.0086
15.17M 0.2735 ± 0.0032 0.4235 ± 0.0022 0.4137 ± 0.0023
Table 19: Model Size Scaling Laws of CodeBrain on the ISRUC_S3 Dataset.
Model Size Cohen’s Kappa Weighted F1 Balanced Accuracy
6.82M 0.7456 ± 0.0083 0.7862 ± 0.0104 0.7514 ± 0.0084
8.49M 0.7486 ± 0.0014 0.7942 ± 0.0040 0.7604 ± 0.0035
10.16M 0.7516 ± 0.0025 0.7985 ± 0.0056 0.7639 ± 0.0079
11.83M 0.7570 ± 0.0082 0.8064 ± 0.0068 0.7734 ± 0.0061
13.50M 0.7620 ± 0.0045 0.8153 ± 0.0091 0.7824 ± 0.0100
15.17M 0.7671 ± 0.0091 0.8202 ± 0.0071 0.7856 ± 0.0031
Figure 16: Model Size Scaling Laws on FACED, SEED-V, and ISRUC_S3.
To better understand the optimization behavior during pretraining, we plot the training loss curves for different model sizes in Figure 17. As expected, larger models consistently achieve lower final training loss, indicating stronger capacity to fit the pretraining objective. The loss reduction is particularly evident when increasing from 3 to 6 layers, while the gain starts to saturate beyond 7 layers. This trend mirrors the downstream performance in Figure 16 and Tables 1719, suggesting that both optimization efficiency and generalization benefit from increased model size—though with diminishing returns as parameter count grows. These results reinforce the scalability of EEGSSM and underscore the importance of balancing capacity with task-specific requirements.
Figure 17: EEGSSM Pre-Training Loss Curve for Different Model Size.
Our results demonstrate consistent improvements in downstream performance as both data volume and model capacity increase, suggesting that EEG foundation models may continue to benefit from further scaling, similar to trends observed in vision and language domains.
Appendix K Low-Resource Comparison With Existing Methods
To evaluate the effectiveness of our model under data-scarce conditions, we conduct experiments where only 30% of the training data is available. Tables 20 and 21 present the results on the FACED (9-class) and SEED-V (5-class) datasets, respectively, comparing our CodeBrain model with several strong baselines. We observe that CodeBrain consistently achieves the best performance among all methods under the 30% setting across all metrics. These improvements are highlighted with underlines in the tables. For reference, we also include the full-data results of CodeBrain. These results highlight CodeBrain’s strong data efficiency and suggest its potential for real-world EEG applications where annotated data is often scarce.
Table 20: Comparison under Partial Data Setting (30%) on the FACED Dataset (9-class).
Methods Cohen’s Kappa Weighted F1 Balanced Accuracy
BIOT (30%) 0.2573 ± 0.0346 0.3501 ± 0.0341 0.3428 ± 0.0329
LaBraM (30%) 0.2672 ± 0.0371 0.3548 ± 0.0325 0.3513 ± 0.0315
CBraMod (30%) 0.3239 ± 0.0265 0.4056 ± 0.0256 0.4035 ± 0.0233
CodeBrain (30%) 0.3356 ± 0.0253 0.4114 ± 0.0225 0.4104 ± 0.0281
CodeBrain (full) 0.5270 ± 0.0039 0.5849 ± 0.0033 0.5816 ± 0.0028
Table 21: Comparison under Partial Data Setting (30%) on the SEED-V Dataset (5-class).
Methods Cohen’s Kappa Weighted F1 Balanced Accuracy
BIOT (30%) 0.1775 ± 0.0425 0.3492 ± 0.0416 0.3505 ± 0.0375
LaBraM (30%) 0.2044 ± 0.0384 0.3700 ± 0.0321 0.3686 ± 0.0305
CBraMod (30%) 0.2291 ± 0.0246 0.3886 ± 0.0255 0.3877 ± 0.0236
CodeBrain (30%) 0.2376 ± 0.0284 0.3943 ± 0.0259 0.3902 ± 0.0271
CodeBrain (full) 0.2735 ± 0.0032 0.4235 ± 0.0022 0.4137 ± 0.0023
Appendix L Limitations and Future Work
Our work presents promising results but also highlights several limitations that offer directions for future exploration. (1) More diverse dataset: Although we pre-train on TUEG, the largest publicly available EEG corpus (over 9000 hours) [66], it has limited channel coverage and originates from a single clinical center, leading to potential distributional bias. This highlights the need for a more diverse, multi-center pre-training dataset with broader channel and task coverage to support robust cross-domain generalization in EEG research. (2) Model scalability: Due to limited data diversity and scale, we used a relatively shallow architecture. While our model design supports deeper networks, effectively training them requires significantly larger and more diverse data. Given the challenges in acquiring large-scale EEG data, future work may involve collaborative efforts to curate richer pre-training corpora that can support scaling to more expressive EEG foundation models. (3) Unimodal focus: Our current framework focuses solely on EEG signals, while other physiological modalities (e.g., EMG, EOG) may also benefit downstream tasks such as emotion recognition and sleep staging. Future work will explore incorporating multimodal EXG signals for more comprehensive representation learning for brain foundation models.
Appendix M Broader impacts
The development of EEG foundation models such as ours marks an important step forward in the intersection of AI and neurophysiological signal processing. Below we outline several potential broader impacts of CodeBrain:
Novel insights for EEG foundation model.
While most existing EEG foundation models rely on Transformer-based architectures, our work demonstrates that non-Transformer designs, particularly those inspired by structured convolutional or state-space models, can achieve competitive performance with lower computational overhead. This may encourage the community to rethink the current architectural paradigm and stimulate research into inductive biases that are better aligned with EEG signal characteristics. We believe this shift could lead to more efficient, interpretable, and scalable EEG modeling frameworks.
Improving interpretability for clinical trust and scientific insight.
Our CodeBrain model enhances interpretability through domain-decoupled tokenization, allowing temporal and frequency components of EEG signals to be disentangled and visualized independently. This design enables a transparent representation of domain-specific brain features (e.g., delta waves in N3 sleep), supporting both clinical trust and neuroscientific interpretation. Such interpretability is critical in high-stakes domains like healthcare, where transparent model decisions facilitate responsible deployment and cross-disciplinary collaboration.
High availability in the real world. By leveraging large-scale pre-training on extensive EEG data, our model can learn general representations applicable to various EEG downstream tasks. This capability is crucial for building BCI systems that can be deployed in different clinical settings, thereby enhancing their accessibility and practicality. Additionally, through improvements in the model architecture, our model achieves acceptable performance while reducing the parameters of the pre-trained model. By decreasing the number of parameters, we lower the memory requirements for deploying the model.

### v2

@@
-CodeBrain: Bridging Decoupled Tokenizer and Multi-Scale Architecture for EEG Foundation Model
+CodeBrain: Towards Decoupled Interpretability and Multi-Scale Architecture for EEG Foundation Model
 Abstract
-Electroencephalography (EEG) provides real-time insights into brain activity and is widely used in neuroscience.
-However, variations in channel configurations, sequence lengths, and task objectives limit the transferability of traditional task-specific models.
-Although recent EEG foundation models (EFMs) aim to learn generalizable representations, they struggle with limited heterogeneous representation capacity and inefficiency in capturing multi-scale brain dependencies.
-To address these challenges, we propose CodeBrain, an efficient EFM structurally aligned with brain organization, trained in two stages.
-(1) We introduce a TFDual-Tokenizer that independently tokenizes heterogeneous temporal and frequency components, enabling a quadratic expansion of the discrete representation space.
-This also offers a degree of interpretability through cross-domain token analysis.
-(2) We propose the EEGSSM, which combines a structured global convolution architecture and a sliding window attention mechanism to jointly model sparse long-range and local dependencies.
-Unlike fully connected Transformer models, EEGSSM better reflects the brain’s small-world topology and efficiently captures EEG’s inherent multi-scale structure.
-EEGSSM is trained with a masked self-supervised learning objective to predict token indices obtained in TFDual-Tokenizer.
-Comprehensive experiments on 10 public EEG datasets demonstrate the generalizability of CodeBrain with linear probing.
-By offering biologically informed and interpretable EEG modeling, CodeBrain lays the foundation for future neuroscience research.
+Electroencephalography (EEG) provides real-time insights into brain activity and supports diverse applications in neuroscience.
+While EEG foundation models (EFMs) have emerged to address the scalability issues of task-specific models, current approaches still yield clinically uninterpretable and weakly discriminative representations, inefficiently capture global dependencies, and neglect important local neural events.
+We present CodeBrain, a two-stage EFM designed to fill this gap.
+In the first stage, we introduce the TFDual-Tokenizer, which decouples heterogeneous temporal and frequency EEG signals into discrete tokens, quadratically expanding the representation space to enhance discriminative power and offering domain-specific interpretability by suggesting potential links to neural events and spectral rhythms.
+In the second stage, we propose the multi-scale EEGSSM architecture, which combines structured global convolution with sliding window attention to efficiently capture both sparse long-range and local dependencies, reflecting the brain’s small-world topology.
+Pretrained on the largest public EEG corpus, CodeBrain achieves strong generalization across 8 downstream tasks and 10 datasets under distribution shifts, supported by comprehensive ablations, scaling-law analyses, and interpretability evaluations.
 Both code and pretraining weights will be released in the future version.
@@
 1 Introduction
-Electroencephalography (EEG) data represent the brain’s electrical activity, captured by placing electrodes on the scalp niedermeyer2005electroencephalography .
-With its high temporal resolution, EEG provides real-time insights into brain dynamics, supporting neuroscience and cognitive research da2013eeg ; binnie2003cognitive .
-To enable automated analysis, researchers have developed various task-specific models, including sleep staging phan2021xsleepnet ; lee2025explainable ; ma2024st , emotion recognition jia2020sst ; guo2024eeg ; wang2024dmmr , motor imagery wang2024st ; ari2024nf ; li2020learning , and other applications guerra2024exploring ; hu2024exploring .
-However, these models are usually trained from scratch and tend to capture task-specific patterns, limiting their applicability beyond the original task.
+Electroencephalography (EEG) captures brain activity via scalp electrodes (niedermeyer2005electroencephalography, ) and provides high temporal-resolution signals for neuroscience and cognitive research (da2013eeg, ).
+To enable automated analysis, researchers have developed various task-specific models for applications such as sleep staging (lee2025explainable, ; ma2025st, ), emotion recognition (jia2020sst, ; liu2024vbh, ), motor imagery (li2020learning, ; jia2020mmcnn, ), and other applications (guerra2024exploring, ; hu2024exploring, ).
+However, building separate models from scratch for each task is resource-intensive and limits scalability, as shared knowledge across tasks cannot be effectively leveraged.
 Moreover, variations in channel configurations and input lengths across EEG tasks further hinder knowledge transfer.
-To tackle these issues, EEG foundation models (EFMs) are developed to learn universal representations for diverse downstream tasks zhou2025brain .
-Inspired by masked self-supervised pretraining in natural language processing van2017neural ; devlin2019bert , current EFMs commonly adopt patch-wise representation learning: EEG signals are divided into fixed-length patches, encoded into latent representations, and trained to reconstruct the masked portions.
-This form of pretraining offers flexibility in downstream applications with varying channel configurations or input lengths by adjusting patch arrangement and number.
-Among reconstruction-based methods, some directly predict the masked raw signals wang2024eegpt ; wang2024cbramod , a task that is challenging due to EEG’s inherent noise and variability.
-Others adopt vector quantization to abstract away signal-level fluctuations jianglarge ; pradeepkumar2025single , and follow a two-stage pipeline: 1) Learn a tokenizer with a codebook to quantize EEG patches into discrete tokens, and 2) pretrain the EFM to predict the token indices of masked patches.
-By acting as a noise-robust latent space, the codebook enables the model to learn more expressive representations.
-Despite these advances, existing EFMs still face challenges.
-Failing to Decouple Heterogeneous EEG Representations.
-Recent EFMs adopt vector quantization with a single codebook for noise-robust representation jianglarge ; pradeepkumar2025single , following image-based VQ-VAE architecture van2017neural , where spatially entangled and homogeneous visual features make a single codebook sufficient zheng2023online ; mentzerfinite .
-However, EEG signals exhibit heterogeneous structures: temporal and frequency components reflect distinct aspects of brain activity miwakeichi2004decomposing .
-Applying a unified codebook to such heterogeneous data can blur domain-specific characteristics, potentially conflating localized temporal events with global frequency patterns liu2024vector and limiting the representation capacity.
-Struggling with Modeling Global Brain Dependencies.
-EEG signals exhibit sparse global dependencies and strong local correlations, reflecting the brain’s small-world topology bullmore2009complex ; bassett2006small ; he2009uncovering .
+To tackle these issues, EEG foundation models (EFMs) are developed to learn universal representations for diverse downstream tasks (zhou2025brain, ).
+Inspired by masked self-supervised pretraining in natural language processing (van2017neural, ; devlin2019bert, ), current EFMs commonly adopt patch-wise representation learning: EEG signals are divided into patches, encoded into latent representations, and trained to reconstruct the masked portions.
+While this approach offers flexibility across varying channel configurations and input lengths by adjusting the patch number and arrangement, direct raw-signal reconstruction (wang2024eegpt, ; wang2024cbramod, ) remains challenging due to the inherent noise and variability of EEG.
+To mitigate this, recent studies have introduced codebook-based tokenization (jianglarge, ; pradeepkumar2025single, ), which abstracts away low-level fluctuations and provides a more robust latent space.
+Despite these advances, existing EFMs still face critical limitations, calling for new architectures.
+Failing to Decouple Heterogeneous EEG for Domain-Specific Interpretability.
+Recent EFMs adopt vector quantization for noise-robust representation (jianglarge, ; pradeepkumar2025single, ), following the VQ-VAE framework originally designed for images, where homogeneous visual features make a single tokenizer sufficient (van2017neural, ; mentzerfinite, ).
+In contrast, EEG exhibits heterogeneous structures: temporal and frequency components reflect distinct aspects of brain activity (miwakeichi2004decomposing, ).
+As illustrated in Fig.
+1(a), signals matched in one domain may diverge in the other.
+Therefore, a mixed tokenizer can conflate domain-specific patterns (liu2024vector, ), weakening representation capacity and producing tokens difficult to align with clinically interpretable neural events or spectral rhythms.
+Struggling with Efficiently Modeling Global Brain Dependencies.
+EEG signals exhibit sparse global dependencies and strong local correlations, reflecting the brain’s small-world topology (bullmore2009complex, ; bassett2006small, ; he2009uncovering, ).
 Efficient modeling of such structure requires capturing relationships in a scalable way.
-However, most EFMs yang2023biot ; jianglarge ; mohammadi2024eeg2rep ; wang2024eegpt ; wang2024cbramod adopt Transformer architectures vaswani2017attention with fully connected self-attention.
-This over-connected design is misaligned with the brain’s sparse structure and struggles to efficiently capture global dependencies due to its quadratic complexity with sequence length tay2021long ; hong2025eegm2 ; tegon2025femba .
+However, most EFMs (yang2023biot, ; jianglarge, ; wang2024eegpt, ; wang2024cbramod, ) adopt Transformer architectures with fully connected self-attention (vaswani2017attention, ).
+This over-connected design is misaligned with the brain’s sparse structure and struggles to efficiently capture global dependencies due to its quadratic complexity with sequence length (tay2021long, ; hong2025eegm2, ; tegon2025femba, ).
 Neglecting Local Dependencies within EEG Patches.
-EEG signals exhibit rich local waveform structures over short temporal windows, reflecting crucial transient neural events tatum2021handbook .
-However, most existing EFMs represent each EEG patch as a single token and apply attention mechanisms only at the patch level wang2024cbramod ; jianglarge , thereby ignoring important local dependencies within patches.
-Figure 1: Overview of our contributions.
-(a) Decoupled vector quantization independently tokenizes temporal and frequency components to preserve heterogeneous EEG structures and enhance representation capacity.
-(b) State Space Model efficiently captures sparse global dependencies across patches.
-(c) Sliding Window Attention models fine-grained local dependencies within patches.
-To address the above challenges, we propose CodeBrain, an efficient EEG foundation model that integrates a decoupled time–frequency tokenizer and a multi-scale architecture.
+EEG signals exhibit rich local waveform structures over short temporal windows, reflecting crucial transient neural events (e.g., sleep waveforms in Fig.
+1(d)) (tatum2021handbook, ).
+However, most existing EFMs represent each EEG patch as a single token and apply attention mechanisms only at the patch level (wang2024cbramod, ; jianglarge, ), thereby ignoring important local dependencies within patches.
+Figure 1: Rationale and overview of CodeBrain beyond existing EFMs.
+(a) EEG signals are heterogeneous, as patches matched in one domain may diverge in the other.
+(b) We then propose a decoupled tokenizer for domain-specific interpretable representations while expanding the representation space.
+(c-d) Inspired by the brain’s small-world topology, a multi-scale architecture further captures inter-patch dependencies efficiently, while modeling overlooked intra-patch neural events.
+(e) These designs deliver performance gains across 10 EEG datasets.
+To address the above challenges, we propose CodeBrain, a novel EEG foundation model that integrates a decoupled tokenizer for domain-specific interpretability with a brain-inspired multi-scale architecture.
 CodeBrain is trained in two stages.
-In the first stage, we introduce TFDual-Tokenizer (Fig.
-1(a)), which independently tokenizes heterogeneous temporal and frequency EEG components.
-To stabilize training, we apply a contrastive learning objective to optimize the temporal codebook.
-In the second stage, we propose EEGSSM, a framework for learning multi-scale EEG representations by masked self-supervised learning.
-EEGSSM predicts token indices obtained from the TFDual-Tokenizer for masked patches.
-To address the inefficiency of over-connected attention and better align with the brain’s small-world structure, we introduce Structured Global Convolution (SGConv) (Fig.
-1(b)) as the backbone.
-Inspired by Structured State Space Models gu2021efficiently ; smith2023simplified ; li2022makes ; gu2022parameterization , SGConv performs efficient global convolution over long EEG sequences, naturally capturing long-range dependencies and local clustering patterns characteristic of brain networks.
-Complementarily, we integrate a Sliding Window Attention (SWA) mechanism (Fig.
-1(c)), adapted from long-sequence modeling beltagy2020longformer , to capture local intra-patch dependencies.
-Together, SGConv and SWA enable EEGSSM to perform efficient and scalable multi-scale EEG representation learning.
-Our detailed contributions can be summarized as follows:
-  • Decoupled Heterogeneous Tokenizer.
-We propose the TFDual-Tokenizer, which independently encodes temporal and frequency components of EEG signals, leading to a quadratic expansion of the discrete representation space.
-It also provides interpretability, supported by visualization analyses in Section 4.5.
-To the best of our knowledge, this is the first vector quantization framework that decouples heterogeneous domains in EEG foundation models.
-  • Efficient Multi-Scale Architecture.
-We design an efficient architecture combining a structured global convolution architecture backbone for global inter-patch dependencies and sliding window attention mechanism for local intra-patch dependencies.
-This design aligns with the brain’s small-world topology, enabling scalable multi-scale representation learning.
-  • Comprehensive Validation.
-We evaluate CodeBrain on 8 diverse downstream brain-computer interface tasks using 10 public EEG datasets.
-Our CodeBrain model demonstrates strong generalization and state-of-the-art performance with linear probing.
+In the first stage, we introduce the TFDual-Tokenizer (Fig.
+1(b)), which decouples temporal and frequency EEG components into discrete tokens to support domain-specific interpretability.
+In the second stage, we develop EEGSSM, a masked self-supervised framework inspired by the brain’s small-world topology.
+EEGSSM adopts a structured global convolution backbone, conceptually related to recent state-space sequence models (smith2023simplified, ; li2022makes, ; gu2022parameterization, ), for sparse and efficient global modeling with a sliding window attention (SWA) mechanism for capturing local neural events overlooked by prior studies(Fig.
+1(c–d)).
+Our detailed contributions are summarized as follows:
+  • Decoupled Tokenizer for Domain-Specific Interpretability.
+We propose the TFDual-Tokenizer, which decouples temporal and frequency EEG components into discrete tokens.
+This design quadratically expands the representation space, and qualitative analyses suggest that the resulting tokens align with neural events and spectral rhythms.
+A contrastive objective is further applied to the temporal branch to stabilize training.
+To the best of our knowledge, this is the first tokenizer in EFMs to provide domain-specific interpretability.
+  • Brain Small-World Topology Inspired Multi-Scale Architecture.
+We design EEGSSM, a patch-wise self-supervised framework for EEG.
+Guided by the brain’s small-world topology, it employs structured global convolution to capture sparse long-range temporal dependencies and sliding window attention to model local neural events.
+In addition, dynamic positional embeddings are used to flexibly learn spatial channel correlations.
+  • Strong Generalization and Comprehensive Validation.
+Pretrained on the largest publicly available EEG corpus, TUEG (obeid2016temple, ), CodeBrain achieves strong performance on 8 downstream tasks across 10 datasets (Fig.
+1(e)) with distribution shifts in cohorts and channel configurations.
+This suggests the model design plays a central role in generalization.
+Comprehensive ablations, scaling-law analyses (3M–150M parameters), together with visualization and quantitative analyses, further confirm its robustness, scalability, and provide domain-specific interpretability.
 2 Related Work
 EEG Foundation Models.
-Inspired by foundation models in vision and language wang2023internimage ; achiam2023gpt , EEG research is shifting from task-specific models jia2020graphsleepnet ; wang2024caresleepnet ; chen2025long to EFMs for learning expressive representations.
+Inspired by foundation models in vision and language (wang2023internimage, ; achiam2023gpt, ), EEG research is shifting from task-specific models (jia2020graphsleepnet, ; wang2024caresleepnet, ; chen2025long, ) to EFMs for learning expressive representations.
 Current EFMs can be divided into two categories.
-1) Contrastive learning-based (CL) chen2025large : BENDER kostas2021bendr first showed the ability of CL for EEG representations.
-Then, the Brant series zhang2023brant ; yuan2024brant ; zhang2024brant enables joint representation learning across physiological signals using CL.
-2) Reconstruction-based: BIOT yang2023biot pioneers cross-modal pretraining for biosignals, including EEG.
-Subsequent models focus specifically on EEG, learning representations by predicting masked discrete tokens jianglarge ; jiangneurolm or reconstructing raw signals wang2024eegpt ; mohammadi2024eeg2rep ; wang2024cbramod .
+1) Contrastive learning-based (CL) (chen2025large, ): BENDER (kostas2021bendr, ) first showed the ability of CL for EEG representations.
+Then, the Brant series (zhang2023brant, ; yuan2024brant, ; zhang2024brant, ) enables joint representation learning across physiological signals using CL.
+2) Reconstruction-based: BIOT (yang2023biot, ) pioneers cross-modal pretraining for biosignals, including EEG.
+Subsequent models focus specifically on EEG, learning representations by predicting masked discrete tokens (jianglarge, ; jiangneurolm, ) or reconstructing raw signals (wang2024eegpt, ; mohammadi2024eeg2rep, ; wang2024cbramod, ).
 Most existing EFMs adopt Transformer architecture, which is suboptimal for EEG due to poor handling of sparse dependencies, quadratic complexity, and ignoring local dependencies by treating each patch as a token.
 EEG Tokenization.
-Tokenization has been key in NLP for generating generalizable and interpretable input representations sennrich2016neural ; kudo2018sentencepiece .
-Inspired by this, early EFMs used patch-based continuous tokenization yang2023biot ; yuan2024brant to handle EEG noise and variability, but without quantization, leading to unbounded and less interpretable representations.
-LaBraM jianglarge introduced vector quantization to learn discrete EEG tokens, following the VQ-VAE design from vision tasks van2017neural .
+Tokenization has been key in NLP for generating generalizable and interpretable input representations (sennrich2016neural, ; kudo2018sentencepiece, ).
+Inspired by this, early EFMs used patch-based continuous tokenization (yang2023biot, ; yuan2024brant, ) to handle EEG noise and variability, but without quantization, leading to unbounded and less interpretable representations.
+LaBraM (jianglarge, ) introduced vector quantization to learn discrete EEG tokens, following the VQ-VAE design from vision tasks (van2017neural, ).
 However, this direct transfer overlooks EEG’s heterogeneous structure, limiting representation capacity.
 Moreover, the tokenizer of LaBraM is trained only on frequency-domain reconstruction due to convergence issues with raw signals reconstruction.
-Later efforts pradeepkumar2025single ; jiangneurolm all relied on a single codebook to encode mixed domain features, limiting both the representation space and interpretability.
-State Space Model Recent works have focused on enhancing classic State-Space Models (SSM) to more efficiently model sequential data using deep learning.
-For example, Syama et al.
-rangapuram2018deep used recurrent neural networks to learn parameters in SSM.
-However, the most significant progress came from Gu et al.
-gu2021efficiently with the introduction of the structured state space model (S4), which reduces the computational complexity of SSM in modeling long sequences using a special state transition matrix gu2020hippo .
+Later efforts (pradeepkumar2025single, ; jiangneurolm, ) all relied on a single codebook to encode mixed domain features, limiting both the representation space and interpretability.
+State Space Model
+Recent works have focused on enhancing classic State-Space Models (SSM) to more efficiently model sequential data using deep learning.
+For example, rangapuram2018deep used recurrent neural networks to learn parameters in SSM.
+However, the most significant progress came from gu2021efficiently with the introduction of the structured state space model (S4), which reduces the computational complexity of SSM in modeling long sequences using a special state transition matrix (gu2020hippo, ).
 These low-rank and normal matrices enable SSM to compute global convolution kernels efficiently through fast Fourier transform across the entire sequence.
-Subsequently, some works have further improved the shortcomings of S4 in areas such as model architecture smith2022simplified and convolution raghu2023sequential , and have started applying it to tasks such as natural language processing dao2022hungry and time series analysis zhou2023deep .
+Subsequently, some works have further improved the shortcomings of S4 in areas such as model architecture (smith2022simplified, ) and convolution (raghu2023sequential, ), and have started applying it to tasks such as natural language processing (dao2022hungry, ) and time series analysis (zhou2023deep, ).
 Recently, some works have also been applied to the EEG data, Tran et al.
-tran2024eeg leveraging SSMs to detect dementia.
+(tran2024eeg, ) leveraging SSMs to detect dementia.
 They extract temporal information from EEG signals through the Mamba architecture and combine it with frequency domain features to better manage the complexity of multivariate EEG.
-In the work of Gui et al.
-gui2024eegmambabidirectionalstatespace , SSM has also become the backbone network of the EEG foundation model.
+In the work of gui2024eegmambabidirectionalstatespace , SSM has also become the backbone network of the EEG foundation model.
 This further highlights the fast reasoning speed and efficient memory usage of the SSM model when processing EEG signals.
 3 Methodology
-3.1 Preliminaries
-Decoupled Tokenization.
-Given a normalized EEG patch $X\in\mathbb{R}^{L}$, where $L$ is the patch length, our goal is to learn a tokenizer that maps $X$ to discrete tokens using two separate codebooks: one for the temporal domain and one for the frequency domain.
-Specifically, let $V^{(t)}\in\mathbb{R}^{K\times D}$ and $V^{(f)}\in\mathbb{R}^{K\times D}$ denote the temporal and frequency codebooks, respectively, where $K$ is the vocabulary size and $D$ is the embedding dimension of each token.
-The tokenizer function is defined as: $z^{(t)},z^{(f)}=f_{\text{tokenizer}}(X)$, where $z^{(t)},z^{(f)}\in\{1,\dots,K\}$ are the indices of the selected codes from the temporal and frequency codebooks.
-EEG Representation Learning.
-Given unlabeled EEG sequences $\mathcal{X}=\{X_{i}\}_{i=1}^{N}$, where each sequence $X_{i}\in\mathbb{R}^{C\times T\times N}$ consists of $C$ channels, a sampling rate $T$, and a duration of $N$ seconds, the goal is to train an encoder $f_{\text{enc}}:\mathbb{R}^{C\times T\times N}\rightarrow\mathbb{R}^{D}$ that produces latent representations $Z_{i}=f_{\text{enc}}(X_{i})$.
 3.1 Model Architecture
-We introduce CodeBrain, a two-stage pretraining framework designed to address the two problem statements described in Section 3.1.
-The model architecture of CodeBrain is illustrated in Fig.
-2.
-In the first stage, we propose TFDual-Tokenizer to decouple temporal and frequency representations of EEG signals, converting the raw signals into noise-robust tokens with potential interpretability.
-In the second stage, we introduce the EEGSSM framework for EEG representation learning.
+We introduce CodeBrain, a two-stage pretraining framework designed to learn interpretable and universal EEG representations (Fig.
+2).
+The model is motivated by complementary goals: 1) domain-specific interpretability via decoupled tokenization of heterogeneous temporal and frequency information, achieved by the proposed TFDual-Tokenizer, and 2) multi-scale modeling of EEG sequences inspired by the brain’s small-world topology, addressed by the EEGSSM framework.
+This design lets Stage 1 learn a tokenizer of patch-level codes, while Stage 2 leverages it for EEG representations.
+We next provide a formal definition of the two stages to clarify their respective roles.
+Stage 1: Decoupled Tokenization.
+Given a normalized EEG patch $\mathbf{x}\in\mathbb{R}^{L}$, where $L$ is the patch length, our goal is to discretize $\mathbf{x}$ into temporal and frequency tokens, enabling domain-specific representation learning.
+Specifically, let $Vt\in\mathbb{R}^{K\times D}$ and $Vf\in\mathbb{R}^{K\times D}$ denote the temporal and frequency codebooks, where $K$ is the vocabulary size and $D$ is the embedding dimension of each token.
+The tokenizer function is defined as: $vt,vf=f_{\text{tokenizer}}(\mathbf{x}),\quad vt,vf\in\mathbb{R}^{D}.$
+Stage 2: EEG Representation Learning.
+Given unlabeled EEG sequences $\mathcal{X}=\{X_{m}\}_{m=1}^{N}$, where each $X_{m}\in\mathbb{R}^{C\times f\times T}$ consists of $C$ channels, sampling rate $f$, and $T$ seconds.
+We divide each sequence into $n$ non-overlapping patches of $t$ seconds, so each patch length is $L=f\cdot t$.
+The goal is to train an encoder $f_{\text{enc}}:\mathbb{R}^{C\times n\times L}\rightarrow\mathbb{R}^{C\times n\times D}$ that produces latent representations $Z_{m}=f_{\text{enc}}(X_{m})$.
 Figure 2: Overview of the CodeBrain framework.
@@
 Each token stream is reconstructed by a decoder to supervise codebook learning in its respective domain.
-Neural Encoder For each patched and normalized EEG input $X\in\mathbb{R}^{C\times T\times N}$, we apply the Discrete Fourier Transform (DFT) cooley1965algorithm to obtain its frequency representation:
-$X[k]=\sum_{n=0}^{N-1}x_{n}\cdot e^{-j\frac{2\pi}{N}kn},\quad k=0,1,\dots,N-1,$ (1)
-where $X[k]$ denotes the $k$-th frequency component, and $n$ denotes the number of patches.
-$X[k]$ is then combined with the raw time-domain signal and fed into the TFConv module, where both are processed in parallel through stacks of convolutional, batch normalization, and ReLU layers.
-The outputs $e_{t}$ and frequency representation $X[K]$ are concatenated to form a time-frequency embedding $e_{p}=Concat\{e_{t},X[K]\}$.
-To obtain patch representation $\tilde{e}_{n}$, we then add a positional embedding $e_{\text{pos}}$ and feed them into TF-Encoder, a Transformer encoder:
-$\tilde{e}_{n}=TransformerEncoder(e_{p}+e_{pos}).$ (2)
-Dual Tokenizer We use two separate tokenizers with distinct codebooks for the temporal and frequency domains, denoted as $v^{(t)}_{j},v^{(f)}_{j}\in\mathbb{R}^{D}$, where $j=1,\dots,K$.
+Neural Encoder For each patch $\mathbf{x}_{i}\in\mathbb{R}^{L}$, we apply the Discrete Fourier Transform (DFT) (cooley1965algorithm, ) to obtain its frequency representation:
+$\mathbf{x}_{i}[k]=DFT(\mathbf{x}_{i})$ (1)
+where $\mathbf{x}_{i}[k]$ denotes the $k$-th frequency component.
+$\mathbf{x}_{i}[k]$ and the $\mathbf{x}_{i}$ are fed into the TFConv module, where they are processed in parallel through stacks of convolutional, batch normalization, and ReLU layers.
+The temporal representation $e^{t}_{i}=TFConv(\mathbf{x}_{i})$ and frequency representation $e_{i}^{f}=TFConv(\mathbf{x}_{i}[k])$ are concatenated to form a time-frequency embedding $e^{p}_{i}=Concat\{e^{t}_{i},e^{f}_{i}\}$.
+To get patch representation $\tilde{e}_{i}$, we then add a positional embedding $e_{\text{pos}}$ and feed it into TF-Encoder:
+$\tilde{e}_{i}=Encoder(e^{p}_{i}+e_{pos}).$ (2)
+We choose a Transformer encoder here since this stage is patch-to-token, where its ability to model local contextual relations makes it well-suited for capturing patch-level patterns.
+To keep the tokenizer channel-agnostic, $e_{\text{pos}}$ is shared temporal embeddings without channel-specific identities.
+Dual Tokenizer We use two separate tokenizers with distinct codebooks for the temporal and frequency domains, denoted as $vt_{j},vf_{j}\in\mathbb{R}^{D}$, where $j=1,\dots,K$.
 Given the patch representation $\tilde{e}_{i}$ from the neural encoder, each tokenizer independently selects the nearest code from its codebook:
-$\hat{z}^{(t)}_{i}=\arg\min_{j}\|\tilde{e}_{i}-v^{(t)}_{j}\|^{2},\quad\hat{z}^{(f)}_{i}=\arg\min_{j}\|\tilde{e}_{i}-v^{(f)}_{j}\|^{2}.$ (3)
-Frequency Codebook Training To train the frequency codebook, we reconstruct two fundamental features—amplitude and phase—from the code embeddings.
-For each EEG patch, we apply the DFT (Eq.
-(1)) to obtain the complex frequency representation: $X[k]=\text{Re}\{X[k]\}+j\cdot\text{Im}\{X[k]\}$
-where $\text{Re}\{X[k]\}$ and $\text{Im}\{X[k]\}$ are the real part and imaginary part respectively, then the amplitude and phase can be calculated as:
-$A_{n}=\sqrt{\text{Re}(X[k])^{2}+\text{Im}(X[k])^{2}},\quad\phi_{n}=\arctan 2\left(\text{Im}(X[k]),\text{Re}(X[k])\right).$ (4)
+$pt_{i}=\arg\min_{j}\|\tilde{e}_{i}-vt_{j}\|^{2},\quad pf_{i}=\arg\min_{j}\|\tilde{e}_{i}-vf_{j}\|^{2},$ (3)
+where $pt_{i}$ and $pf$ denote the closest positions for the embeddings in the temporal and frequency domain codebook. The effectiveness of the Dual Tokenizer is based on the following proposition:
+Proposition 2.1 Decoupling temporal and frequency codebooks yields representations that are no less effective than those from a joint codebook.
+Proof. See Appendix C.
+For this Proposition, we provide empirical validation in Sections 4.4, 4.5 and analysis of Dual Tokenizer ’s interpretable structure in Appendix A.
+Frequency Codebook Training To train the frequency codebook, we reconstruct amplitude and phase from the code embeddings.
+For each EEG patch, we apply the DFT to obtain the frequency representation: $\mathbf{x}_{i}[k]=\text{Re}\{\mathbf{x}_{i}[k]\}+j\cdot\text{Im}\{\mathbf{x}_{i}[k]\}$ where $\text{Re}\{\mathbf{x}_{i}[k]\}$ and $\text{Im}\{\mathbf{x}_{i}[k]\}$ are the real part and imaginary part respectively, then the amplitude and phase can be calculated as:
+$A_{i}=\sqrt{\text{Re}(\mathbf{x}_{i}[k])^{2}+\text{Im}(\mathbf{x}_{i}[k])^{2}},\quad\phi_{i}=\arctan 2\left(\text{Im}(\mathbf{x}_{i}[k]),\text{Re}(\mathbf{x}_{i}[k])\right).$ (4)
 We use z-score normalization to ensure stable training.
-The code embedding $v^{(f)}_{\hat{z}_{n}}$, retrieved from the frequency codebook, is passed through the F-Decoder, which consists of a Transformer encoder followed by two linear layers:
-$y_{n}^{A}=TF^{(f)}(MLP(\tilde{e}_{n})),\quad y_{n}^{P}=TF^{(f)}(MLP(\tilde{e}_{n})).$ (5)
-where $y_{n}^{A}$ and $y_{n}^{P}$ are the predicted amplitude and phase, respectively.
-$TF^{(f)}$ denotes a Transformer encoder.
-The frequency codebook’s training objective frequency codebook is the mean squared error (MSE) loss:
-$\mathcal{L}_{n}^{f}=||y_{n}^{A}-A_{n}||^{2}_{2}+||y_{n}^{P}-\phi_{n}||^{2}_{2}.$ (6)
-Temporal Codebook Training Direct reconstruction of temporal features might lead to non-convergence jianglarge .
+The code embedding ${vf}_{i}$, retrieved from the frequency codebook, is passed through the F-Decoder, which consists of a Transformer encoder followed by two linear layers:
+$y_{i}^{A}=Encoder(MLP(\tilde{e}_{i})),\quad y_{i}^{P}=Encoder(MLP(\tilde{e}_{i})).$ (5)
+where $y_{i}^{A}$ and $y_{i}^{P}$ are the predicted amplitude and phase, respectively. The frequency codebook’s training objective is the mean squared error (MSE) loss:
+[…]
+Temporal Codebook Training
+Direct reconstruction of temporal features might lead to non-convergence (jianglarge, ).
 To address this, we combine contrastive loss with reconstruction loss to train a temporal codebook.
-Inspired by studies on physiological signals kiyasseh2021clocs , we assume temporal dependencies exist between different EEG patches, especially within the same channel.
-For an EEG sample $X\in\mathbb{R}^{C\times T\times N}$, we split it into two halves of length $N/2$.
+Inspired by studies on physiological signals (kiyasseh2021clocs, ), we assume temporal dependencies exist between EEG segments, especially within the same channel.
+For an EEG segment $X_{m}\in\mathbb{R}^{C\times n\times L}$, we split it into two halves of length $n/2$.
 We use TF-Encoder in Eq.
-(2) to obtain the representations of these two segments $X_{h1},X_{h2}\in\mathbb{R}^{C\times T\times N/2}$ separately:
-$h_{n}=TransformerEncoder(X_{{n}}),X_{n}\in\mathbb{R}^{C\times(T\times N/2)}.$ (7)
-We encourage the latent representations of different segments within a single sample $h_{n}$ and $\tilde{h}_{n}$ to be similar, while making the latent representations of the same segment across different samples $h_{n}$ and $h_{k}$ as distinct as possible.
-We use SimCLR loss chen2020simple to train the temporal codebook:
-$\mathcal{L}_{CL}=-\log\frac{\exp\left(\mathrm{sim}({h}_{n},\tilde{h}_{n})/\tau\right)}{\sum\limits_{k=1}^{2N}\mathbb{1}_{[k\neq i]}\exp\left(\mathrm{sim}({h}_{n},{h}_{k})/\tau\right)}\quad,$ (8)
-where $\mathrm{sim}(\mathbf{h}_{i},\mathbf{h}_{j})$ is the cosine similarity, $\tau>0$ is the temperature parameter, and $\mathbb{1}_{[k\neq i]}$ is an indicator function used to exclude itself.
-We introduce a T-Decoder to reconstruct raw signals from the temporal code embedding $\tilde{e}_{n}$, which consists of a Transformer encoder $TF^{(t)}$ followed by a linear projection.
+(2) to obtain the representations of these two parts $X_{m1},X_{m2}\in\mathbb{R}^{C\times\frac{2}{n}\times L}$ separately:
+$e^{h}_{mi}=Encoder(X_{mi}),i\in\{1,2\}.$ (7)
+We encourage the latent representations of different parts within a single segment $e^{h}_{m1}$ and $e^{h}_{m2}$ to be similar, while making those of the same part across different segments $e^{h}_{mi}$ and $e^{h}_{sk}$ as distinct as possible, where $X_{s}\in\mathbb{R}^{C\times n\times L},s\neq m,k\neq i$.
+SimCLR loss (chen2020simple, ) is used for training:
+$\mathcal{L}^{CL}_{m}=-\log\frac{\exp\left(\mathrm{sim}(e^{h}_{m1},e^{h}_{m2})/\tau\right)}{\sum\limits_{k=1}^{2N}\mathbb{1}_{[k\neq i]}\exp\left(\mathrm{sim}(e^{h}_{mi},e^{h}_{sk})/\tau\right)},$ (8)
+where $\mathrm{sim}(e^{h}_{m1},e^{h}_{m2})$ is the cosine similarity, $\tau>0$ is the temperature parameter, and $\mathbb{1}_{[k\neq i]}$ is an indicator function used to exclude itself.
+We introduce a T-Decoder to reconstruct raw signals from the temporal code embedding $\tilde{e}_{i}$, which consists of a Transformer encoder followed by a linear projection.
 Let $y_{t}$ be the Transformer output from Eq.
 (2); the overall training objective is:
-$\mathcal{L}_{t}=\mathcal{L}_{CL}+||y_{t}-x_{t}||^{2}_{2},\quad y_{t}=TF^{(t)}(MLP(\tilde{e}_{n})).$ (9)
-Finally, the training objective for the TFDual-Tokenizer is defined as:
-$\mathcal{L}_{codebook}=\sum_{x\in\mathcal{D}}\sum_{i}\underbrace{||y_{n}^{A}-A_{n}||^{2}_{2}+||y_{n}^{P}-\phi_{n}||^{2}_{2}}_{\text{reconstruction loss}}+\underbrace{\|sg(\tilde{e}_{n})-v_{j}^{(t)}\|^{2}+\|sg(\tilde{e}_{n})-v_{j}^{(f)}\|^{2}}_{\text{codebook loss}}+\mathcal{L}_{t},$ (10)
-where $sg(\cdot)$ denotes stop-gradient, avoiding the update of encoder parameters.
+[…]
+Finally, the training objective for the TFDual-Tokenizer can be defined as:
+$\mathcal{L}_{\text{tokenizer}}=\sum_{X_{m}\in\mathcal{X}}\mathcal{L}^{CL}_{m}+\sum_{X_{m}\in\mathcal{X}}\sum_{i=0}^{n}\ \mathcal{L}_{i}^{f}+\mathcal{L}^{t}_{i}$ $+\underbrace{\|\mathrm{sg}(\tilde{e}_{i})-vt_{pt_{i}}\|^{2}+\|\mathrm{sg}(\tilde{e}_{i})-vf_{pf_{i}}\|^{2}}_{\text{codebook loss}}$ (10)
+$+\underbrace{\|\tilde{e}_{i}-\mathrm{sg}(vt_{pt_{i}})\|^{2}+\|\tilde{e}_{i}-\mathrm{sg}(vf_{pf_{i}})\|^{2}}_{\text{commitment loss}},$
+where $\mathrm{sg}(\cdot)$ denotes stop-gradient to avoid updating encoder parameters.
 3.3 EEGSSM Pretraining
 In this stage, we introduce a novel convolutional structured state space model framework, called EEGSSM, composed of multiple EEGSSM blocks.
-First, raw EEG signals are projected into a higher-dimensional embedding space via a 1D convolutional layer.
+To adapt to unseen channels, we first learn dynamic positional embeddings that capture inter-channel relationships through a lightweight convolutional module.
 The resulting features are processed by EEGSSM blocks.
 Another 1D convolutional layer then maps the output back to the token space for reconstructing the indices of masked tokens produced by the TFDual-Tokenizer.
-EEGSSM Block Our EEGSSM block is composed of several blocks, which are integrated together through a residual connection mechanism.
-An EEGSSM block consists of an RMS Norm layer, SGConv layer, SWA layer, and a gating component.
-Root Mean Square Layer Normalization (RMSNorm) zhang2019rootmeansquarelayer can be considered an improved version of LayerNorm, normalizing by using the square root of the mean to reduce the impact of noise.
+EEGSSM Block
+Our EEGSSM block is composed of several blocks, which are integrated together through a residual connection mechanism.
+An EEGSSM block consists of a Layer Normalization, SGConv layer, SWA layer, and a gating component.
 Afterward, we feed the intermediate variables into the SGConv layer to obtain a global receptive field through convolution SSM.
-SGConv Layer. SGConv is a structured SSM model using convolution architecture (See supplemental material), and its convolution structure can be represented as a FFT formula:
+SGConv Layer. SGConv is a structured SSM model (see Appendix B) using convolution architecture, and its convolution structure can be represented as a DFT formula:
 $y=F^{-1}_{N}D_{k}F_{N}u,D_{k}=\text{diag}(\overline{K}F_{N}),$ (11)
-where $F_{N}$ denotes the DFT matrix of size $N$.
-This FFT convolution has a computational complexity of $O(nlog(n))$.
+where $F_{N}$ denotes the DFT matrix of size $N$, and the convolution can be computed in $O(N\log N)$ via FFT.
 As a type of convolutional SSM, SGConv improves the convolution kernel $\overline{K}$ in Eq.
@@
 $\overline{K}=\frac{1}{Z}[k_{0},k_{1},...,k_{N-1}],k_{i}=\alpha^{i}\mathrm{Upsample}_{2^{max[i-1,0]}d}(w_{i}),$ (12)
-where $\alpha$ denotes the decay coefficient, usually chosen to be 1/2, inducing the decaying structure, and $\text{Upsample}(x)$ denotes upsample $x$ to length $l$.
+where $\alpha$ denotes the decay coefficient, usually chosen to be 0.5, inducing the decaying structure, and $\text{Upsample}(x)$ denotes upsample $x$ to length $l$.
 We also introduce the normalization constant $Z$ to ensure that the convolution operation does not change the scale of the input.
@@
 Furthermore, SWA maintains linear computational complexity to ensure training speed remains largely unaffected.
-At the end of the EEGSSM block, we use a gating mechanism to control the output of the block.
-We employ a gated unit similar to Wavenet obeid2016temple , which can suppress useless or irrelevant features and help stabilize training in deep networks.
+Gate Mechanism.
+we use a gating mechanism to control the output of the block.
+We employ a gated unit similar to Wavenet (obeid2016temple, ), which can suppress useless or irrelevant features and help stabilize training in deep networks.
 We concatenate the global features output $y_{sg}$ by the SGConv layer with the local features output $y_{swa}$ by SWA and feed them into a gated unit:
@@
 $y_{1}=\mathrm{Conv}(z),y_{2}=\mathrm{Conv}(z)$ (14)
-where $tanh(\cdot)$ and $\sigma(\cdot)$ are the tanh function and sigmoid function, $\odot$ denotes an element-wise multiplication operator, $W_{f}$ and $W_{g}$ are learnable convolution filters.
+where $tanh(\cdot)$ and $\sigma(\cdot)$ are tanh function and sigmoid function, $\odot$ denotes an element-wise multiplication operator, $W_{f}$ and $W_{g}$ are learnable convolution filters.
 $y_{1}$ becomes the input to the next block, while $y_{2}$ will be aggregated to the output of SSM blocks through a skip connection.
-Pre-Training Objective To help the EEGSSM model learn general EEG representations, we use a Masked Autoencoder (MAE) for self-supervised pre-training.
-For a patched sample $X=\{x_{i}\mid i\in[1,2,\ldots,C]\}$, we randomly generate a mask =$\mathcal{M}=\{m_{i}\mid i\in[1,2,\ldots,C]\}$ from a Bernoulli distribution of $r$ proportion, where $m_{ij}\in\{0,1\}$.
+Pre-Training Objective
+To help the EEGSSM model learn general EEG representations, we use a Masked Autoencoder (MAE) for self-supervised pre-training.
+For a patched sample $X=\{x_{i}\mid i\in[1,2,\ldots,C]\}$, we randomly generate a mask $\mathcal{M}=\{m_{i}\mid i\in[1,2,\ldots,C]\}$ from a Bernoulli distribution of $r$ proportion, where $m_{i}\in\{0,1\}$.
 We reconstruct the token indices of masked EEG patches from the TFDual-Tokenizer by cross-entropy loss.
-Let $y_{i}$ denote the output of the EEGSSM block, the probability that the EEG signal matches the corresponding token $v_{i}$ in the codebooks:
+Let $y_{i}$ denotes the output of the EEGSSM block, the probability that the EEG signal matches the corresponding token $v_{i}$ in the codebooks:
 $p(v_{i}|x_{i})=softmax(Conv1D(y_{i})).$ (15)
@@
 4.1 Datasets
-Pre-Training We pretrain CodeBrain on the TUH EEG Corpus (v2.0.1, modified in March 2025) obeid2016temple , the largest publicly available EEG dataset to date.
+Pre-Training.
+We pretrain CodeBrain on the TUH EEG Corpus (obeid2016temple, ), the largest publicly available EEG dataset to date.
 Data processing follows a standardized pipeline: recordings shorter than 5 minutes are excluded, and the first and last minute of each segment is removed.
-We retain 19 commonly used EEG channels (C3, C4, Cz, F3, F4, Fp1, Fp2, F7, F8, Fz, O1, O2, P3, P4, Pz, T3, T4, T5, T6), selected based on the international 10–20 system for electrode placement hh1958ten , apply band-pass filtering (0.3–75 Hz), and notch filtering at 60 Hz to remove noise.
+We retain 19 commonly used EEG channels (C3, C4, Cz, F3, F4, Fp1, Fp2, F7, F8, Fz, O1, O2, P3, P4, Pz, T3, T4, T5, T6), selected based on the international 10–20 system for electrode placement (hh1958ten, ).
+We apply band-pass filtering (0.3–75 Hz), and notch filtering at 60 Hz to remove noise.
 The data is resampled to 200 Hz and divided into 30-second non-overlapping segments.
-Segments with absolute amplitudes over $100\leavevmode\nobreak\ \mu\mathrm{V}$ are filtered out.
+Segments with absolute amplitudes over $100\penalty 10000\ \mu\mathrm{V}$ are filtered out.
 To normalize the signals, each value is divided by 100.
@@
 Downstream Tasks.
-We evaluate CodeBrain on 8 downstream tasks across 10 publicly available EEG datasets, covering a diverse range of applications.
+We evaluate CodeBrain on 8 downstream tasks across 10 public EEG datasets, which span diverse applications and exhibit distribution shifts from the pretraining dataset, to assess generalizability.
 Detailed dataset configurations are in Table 1.
 We perform cross-subject or cross-session splits with strict separation between training, validation, and test sets.
-For FACED chen2023large , we use 80 subjects for training, 20 for validation, and the remaining 23 for testing.
-In SEED-V liu2021comparing , we perform a leave-one-session-out manner.
-ISRUC_S3 khalighi2016isruc consists of 10 subjects, for which we apply an 8:1:1 cross-subject split.
+For FACED, we use 80 subjects for training, 20 for validation, and the remaining 23 for testing.
+In SEED-V, each session consists of 15 trials, which are evenly divided into training, validation, and test sets.
+ISRUC_S3 consists of 10 subjects, for which we apply an 8:1:1 cross-subject split.
 MentalArithmetic consists of 36 subjects, and we use a 7:1:1 cross-subject split.
-BCIC2020-T3 jeong20222020 follows the official competition protocol.
-In CHB-MIT shoeb2009application , we use recordings from 19 subjects for training, and from 2 subjects each for validation and testing.
-Additional dataset details are provided in the supplementary material.
+BCIC2020-T3 follows the official competition protocol.
+In CHB-MIT, we use recordings from 19 subjects for training, and from 2 subjects each for validation and testing.
+Additional dataset details are provided in the Appendix F.
 Table 1: Summary of Downstream Tasks and Associated EEG Datasets.
-Downstream Tasks Datasets Shape #Samples Class
-Emotion Recognition FACED chen2023large 32 channels * 10s 10,332 9-class
-SEED-V liu2021comparing 62 channels * 1s 117,744 5-class
-Sleep Staging ISRUC_S1 khalighi2016isruc 6 channels * 30s 86,320 5-class
-ISRUC_S3 khalighi2016isruc 6 channels * 30s 8,500 5-class
-Imagined Speech Classification BCIC 2020-T3 jeong20222020 64 channels * 3s 6,000 5-class
-Mental Stress Detection Mental Arithmetic mumtaz2016eeg 20 channels * 5s 1,707 2-class
-Seizure Detection CHB-MIT shoeb2009application 16 channels * 10s 326,993 2-class
-Motor Imagery Classification SHU-MI goldberger2000physiobank 32 channels * 4s 11,988 2-class
-Event Type Classification TUEV obeid2016temple 16 channels * 5s 112,491 6-class
-Abnormal Detection TUAB obeid2016temple 16 channels * 10s 409,455 2-class
+Downstream Tasks Datasets #Channels Length #Samples Class
+Emotion Recognition FACED (chen2023large, ) 32 10s 10,332 9-class
+SEED-V (liu2021comparing, ) 62 1s 117,744 5-class
+Sleep Staging ISRUC_S1 (khalighi2016isruc, ) 6 30s 86,320 5-class
+ISRUC_S3 (khalighi2016isruc, ) 6 30s 8,500 5-class
+Imagined Speech Classification BCIC 2020-T3 (jeong20222020, ) 64 3s 6,000 5-class
+Mental Stress Detection Mental Arithmetic (mumtaz2016eeg, ) 20 5s 1,707 2-class
+Seizure Detection CHB-MIT (shoeb2009application, ) 16 10s 326,993 2-class
+Motor Imagery Classification SHU-MI (goldberger2000physiobank, ) 32 4s 11,988 2-class
+Event Type Classification TUEV (obeid2016temple, ) 16 5s 112,491 6-class
+Abnormal Detection TUAB (obeid2016temple, ) 16 10s 409,455 2-class
+Table 2: Comparison results of different methods on downstream tasks.
+Methods FACED (9-Class) SEED-V (5-Class)
+Cohen’s Kappa Weighted F1 Balanced Acc Cohen’s Kappa Weighted F1 Balanced Acc
+BENDR 0.4716 ± 0.0095 0.5340 ± 0.0086 0.5320 ± 0.0083 0.0335 ± 0.0062 0.2026 ± 0.0330 0.2231 ± 0.0059
+BIOT 0.4476 ± 0.0254 0.5136 ± 0.0112 0.5118 ± 0.0118 0.2261 ± 0.0262 0.3856 ± 0.0203 0.3837 ± 0.0187
+LaBraM 0.4698 ± 0.0102 0.5288 ± 0.0188 0.5273 ± 0.0107 0.2386 ± 0.0209 0.3974 ± 0.0111 0.3976 ± 0.0138
+EEGPT 0.4639 ± 0.0023 0.3924 ± 0.0017 0.4607 ± 0.0014 0.1323 ± 0.0062 0.3090 ± 0.0052 0.3061 ± 0.0044
+CBraMod 0.5041 ± 0.0122 0.5618 ± 0.0093 0.5509 ± 0.0089 0.2569 ± 0.0143 0.4101 ± 0.0108 0.4091 ± 0.0097
+CodeBrain 0.5406 ± 0.0084 0.5953 ±0.0113 0.5941 ± 0.0098 0.2735 ± 0.0032 0.4235 ± 0.0022 0.4137 ± 0.0023
+Methods ISRUC_S3 (5-Class) BCIC 2020-T3 (5-Class)
+Cohen’s Kappa Weighted F1 Balanced Acc Cohen’s Kappa Weighted F1 Balanced Acc
+BENDR 0.5995 ± 0.0151 0.6789 ± 0.0142 0.6352 ± 0.0095 0.0607 ± 0.0093 0.2379 ± 0.0165 0.2485 ± 0.0075
+BIOT 0.7168 ± 0.0119 0.7834 ± 0.0096 0.7598 ± 0.0109 0.3650 ± 0.0176 0.4917 ± 0.0079 0.4920 ± 0.0086
+LaBraM 0.7194 ± 0.0162 0.7843 ± 0.0189 0.7617 ± 0.0122 0.3800 ± 0.0242 0.5054 ± 0.0205 0.5060 ± 0.0155
+EEGPT 0.6160± 0.0856 0.6375 ± 0.0632 0.6650 ± 0.0311 0.0567 ± 0.0164 0.2441 ± 0.0105 0.2453 ± 0.0131
+CBraMod 0.7407 ± 0.0251 0.8056 ± 0.0219 0.7844 ± 0.0126 0.4216 ± 0.0163 0.5383 ± 0.0096 0.5373 ± 0.0108
+CodeBrain 0.7671 ± 0.0091 0.8202 ± 0.0071 0.7856 ± 0.0031 0.5127 ± 0.0065 0.6101 ± 0.0053 0.6101 ± 0.0052
+Methods Mental Arithmetic (2-Class) CHB_MIT (2-Class)
+AUROC AUC-PR Balanced Acc AUROC AUC-PR Balanced Acc
+BENDR 0.6248 ± 0.0765 0.3661 ± 0.0672 0.5681 ± 0.0448 0.8632 ± 0.0526 0.3071 ± 0.1240 0.5609 ± 0.0432
+BIOT 0.7536 ± 0.0144 0.6004 ± 0.0195 0.6875 ± 0.0186 0.8761 ± 0.0284 0.3277 ± 0.0460 0.7068 ± 0.0457
+LaBraM 0.7721 ± 0.0093 0.5999 ± 0.0155 0.6909 ± 0.0125 0.8679 ± 0.0199 0.3287 ± 0.0402 0.7075 ± 0.0358
+EEGPT 0.7162 ± 0.0171 0.5081 ± 0.0275 0.5597 ± 0.0171 0.8892 ± 0.0066 0.3073 ± 0.0641 0.5481 ± 0.0151
+CBraMod 0.7905 ± 0.0073 0.6267 ± 0.0099 0.7256 ± 0.0132 0.8892 ± 0.0154 0.3689 ± 0.0382 0.7398 ± 0.0284
+CodeBrain 0.8707 ± 0.0209 0.7177 ± 0.0421 0.7514 ± 0.0203 0.8961 ± 0.0174 0.4377 ± 0.0288 0.7273 ± 0.0240
 4.2 Experiment Settings
-Two-Stage Training Setup.
+Experiment Setup.
+(1) Pretraining Setup.
 All experiments are conducted on NVIDIA 40GB A100 GPUs.
-In Stage 1, the TFDual-Tokenizer is trained with temporal and frequency codebooks of 4096 codes (32 dimensions) for 20 epochs, using a batch size of 256 and a learning rate of 1e-4, across six A100 GPUs for approximately ten hours.
-In Stage 2, an 8-layer EEGSSM backbone with a masking ratio of 0.5 is trained for 10 epochs, using a batch size of 256 on two A100 GPUs for about one day.
-Fintuning Strategy.
-We adopt a linear probing setup to evaluate the quality of pretrained representations.
+The TFDual-Tokenizer is trained with temporal and frequency codebooks of 4096 codes (32 dimensions) for 20 epochs, and an 8-layer EEGSSM backbone with a masking ratio of 0.5 is trained for 10 epochs.
+(2) Finetuning Strategy.
+We evaluate the quality of the pretrained representations under full finetuning.
 All downstream task datasets are resampled to 200 Hz to match the pretraining configuration.
-A three-layer MLP is applied.
-The first layer aggregates information across all EEG channels.
-The second layer compresses the entire $x$-second sequence (depending on task configuration) into a 200-dimensional vector.
-The final layer maps this representation to the target class.
-Each hidden layer uses ELU activation and dropout to improve generalization.
+A three-layer MLP is applied to aggregate channel information, compress the $x$-second sequence, and map the representation to the target class, with activation and dropout between layers.
 Baselines.
 We compare our model with five publicly available EFMs that have released pretrained weights, representing different pretraining strategies.
-BENDR kostas2021bendr adopts contrastive learning to learn general EEG representations.
-BIOT yang2023biot uses patch-based continuous tokenization.
-LaBraM introduces discrete neural tokens through vector quantization jianglarge .
-EEGPT wang2024eegpt and CBraMod wang2024cbramod both rely on masked reconstruction of raw EEG signals.
+BENDR (kostas2021bendr, ) adopts contrastive learning framework.
+BIOT (yang2023biot, ) uses patch-based continuous tokenization.
+LaBraM introduces discrete neural tokens through vector quantization (jianglarge, ).
+EEGPT (wang2024eegpt, ) and CBraMod (wang2024cbramod, ) rely on masked reconstruction of raw EEG signals.
 Evaluation Metric.
@@
 All experiments are repeated with five random seeds, and we report the mean and standard deviation.
+More details about hyperparameter, baseline, and evaluation metric are provided in Appendix H, G.
 4.3 Comparison with Baselines
 We ensure consistent data splits across all baselines.
-Results are reported on six representative downstream datasets, with additional results provided in the supplementary material.
+Results are reported on six representative downstream datasets, with additional results provided in Appendix I.
 As shown in Table 2, CodeBrain achieves consistent performance gains compared to baselines.
-For multi-class classification, it achieves the largest gain of +0.0911 in Cohen’s Kappa (21.6%), +0.0718 in Weighted F1 score (13.3%), +0.0728 in Balanced Acc (13.5%) on BCIC 2020-T3 over the strongest baseline wang2024cbramod .
-For binary classification, it achieves the largest gain of +0.0802 in AUROC (10.1%), +0.0910 in AUC-PR (14.5%) and +0.0258 in Balanced Acc (3.6%) on Mental Arithmetic over the strongest baseline wang2024cbramod .
+For multi-class classification, it achieves the largest gain of +0.0911 in Cohen’s Kappa (21.6%), +0.0718 in Weighted F1 score (13.3%), +0.0728 in Balanced Acc (13.5%) on BCIC 2020-T3 over the strongest baseline (wang2024cbramod, ).
+For binary classification, it achieves the largest gain of +0.0802 in AUROC (10.1%), +0.0910 in AUC-PR (14.5%) and +0.0258 in Balanced Acc (3.6%) on Mental Arithmetic over the strongest baseline.
 These results demonstrate the superior generalizability of CodeBrain.
-Table 2: Comparison results of different methods on downstream tasks.
-Methods FACED (9-Class) SEED-V (5-Class)
-Kappa W-F1 Bal. Acc Kappa W-F1 Bal. Acc
-BENDRkostas2021bendr 0.4716 ± 0.0095 0.5340 ± 0.0086 0.5320 ± 0.0083 0.0335 ± 0.0062 0.2026 ± 0.0330 0.2231 ± 0.0059
-BIOTyang2023biot 0.4476 ± 0.0254 0.5136 ± 0.0112 0.5118 ± 0.0118 0.2261 ± 0.0262 0.3856 ± 0.0203 0.3837 ± 0.0187
-LaBraMjianglarge 0.4698 ± 0.0102 0.5288 ± 0.0188 0.5273 ± 0.0107 0.2386 ± 0.0209 0.3974 ± 0.0111 0.3976 ± 0.0138
-EEGPTwang2024eegpt 0.4639 ± 0.0023 0.3924 ± 0.0017 0.4607 ± 0.0014 0.1323 ± 0.0062 0.3090 ± 0.0052 0.3061 ± 0.0044
-CBraModwang2024cbramod 0.5041 ± 0.0122 0.5618 ± 0.0093 0.5509 ± 0.0089 0.2569 ± 0.0143 0.4101 ± 0.0108 0.4091 ± 0.0097
-CodeBrain 0.5270 ± 0.0039 0.5849 ± 0.0033 0.5816 ± 0.0028 0.2735 ± 0.0032 0.4235 ± 0.0022 0.4137 ± 0.0023
-Methods ISRUC_S3 (5-Class) BCIC 2020-T3 (5-Class)
-Kappa W-F1 Bal. Acc Kappa W-F1 Bal. Acc
-BENDRkostas2021bendr 0.5995 ± 0.0151 0.6789 ± 0.0142 0.6352 ± 0.0095 0.0607 ± 0.0093 0.2379 ± 0.0165 0.2485 ± 0.0075
-BIOTyang2023biot 0.7168 ± 0.0119 0.7834 ± 0.0096 0.7598 ± 0.0109 0.3650 ± 0.0176 0.4917 ± 0.0079 0.4920 ± 0.0086
-LaBraMjianglarge 0.7194 ± 0.0162 0.7843 ± 0.0189 0.7617 ± 0.0122 0.3800 ± 0.0242 0.5054 ± 0.0205 0.5060 ± 0.0155
-EEGPTwang2024eegpt 0.6160± 0.0856 0.6375 ± 0.0632 0.6650 ± 0.0311 0.0567 ± 0.0164 0.2441 ± 0.0105 0.2453 ± 0.0131
-CBraModwang2024cbramod 0.7407 ± 0.0251 0.8056 ± 0.0219 0.7844 ± 0.0126 0.4216 ± 0.0163 0.5383 ± 0.0096 0.5373 ± 0.0108
-CodeBrain 0.7671 ± 0.0091 0.8202 ± 0.0071 0.7856 ± 0.0031 0.5127 ± 0.0065 0.6101 ± 0.0053 0.6101 ± 0.0052
-Methods Mental Arithmetic (2-Class) CHB_MIT (2-Class)
-AUROC AUC-PR Bal. Acc AUROC AUC-PR Bal. Acc
-BENDRkostas2021bendr 0.6248 ± 0.0765 0.3661 ± 0.0672 0.5681 ± 0.0448 0.8632 ± 0.0526 0.3071 ± 0.1240 0.5609 ± 0.0432
-BIOTyang2023biot 0.7536 ± 0.0144 0.6004 ± 0.0195 0.6875 ± 0.0186 0.8761 ± 0.0284 0.3277 ± 0.0460 0.7068 ± 0.0457
-LaBraMjianglarge 0.7721 ± 0.0093 0.5999 ± 0.0155 0.6909 ± 0.0125 0.8679 ± 0.0199 0.3287 ± 0.0402 0.7075 ± 0.0358
-EEGPTwang2024eegpt 0.7162 ± 0.0171 0.5081 ± 0.0275 0.5597 ± 0.0171 0.8892 ± 0.0066 0.3073 ± 0.0641 0.5481 ± 0.0151
-CBraModwang2024cbramod 0.7905 ± 0.0073 0.6267 ± 0.0099 0.7256 ± 0.0132 0.8892 ± 0.0154 0.3689 ± 0.0382 0.7398 ± 0.0284
-CodeBrain 0.8707 ± 0.0209 0.7177 ± 0.0421 0.7514 ± 0.0203 0.8961 ± 0.0174 0.4377 ± 0.0288 0.7273 ± 0.0240
-4.4 Ablation Study
-We conduct ablation studies on three datasets using the same five seeds as in the main experiments to evaluate the key components in CodeBrain, as shown in Table 3.
-Additional results on mask ratio and architecture efficiency are in the supplementary material.
-Below is the detailed analysis:
-(1) Tokenizer configuration: we compare the proposed TFDual-Tokenizer (Dual) with variants using only one domain codebook (Temporal or Frequency), or a shared codebook jointly reconstructing both domains (Mixed).
-Across all datasets, the Dual codebook consistently yields superior performance, suggesting that maintaining separate temporal and frequency codebooks captures complementary information more effectively than merging them.
-(2) Sliding Window Attention (SWA): Including the SWA module consistently improves performance, confirming its regularization effect in capturing local dependencies.
-Notably, on the SEED-V dataset with 1s samples, SWA yields the largest gain, confirming its effectiveness for short-duration inputs.
-(3) Contrastive learning (CL): We evaluate the impact of contrastive learning in TFDual-Tokenizer pretraining. It also leads to consistent gains, indicating a better capture of temporal patterns.
-(4) Scaling Laws: Prior work wang2024cbramod ; jianglarge has explored scaling with 1 to 1000 hours of EEG data for pretraining and model sizes up to 5M parameters.
-We extend this by analyzing performance using 1k to 9k hours of pretraining data and larger models.
-As shown in Figure 3, Kappa, our primary evaluation metric, consistently improves with more data and parameters, suggesting that further scaling may lead to continued performance gains.
 Table 3: The results of ablation studies for tokenizer configurations and module components.
-Dataset Codebook SWA CL Cohen’s Kappa Weighted F1 Balanced Accuracy
-FACED 9-Class Dual 0.5270 $\pm$ 0.0039 0.5849 $\pm$ 0.0033 0.5816 $\pm$ 0.0028
+Dataset Codebook CL SWA SGConv Gate Cohen’s Kappa Weighted F1 Balanced Accuracy
+FACED 9-Class Dual 0.5406 $\pm$ 0.0084 0.5953 $\pm$ 0.0113 0.5941 $\pm$ 0.0098
 Temporal 0.4618 $\pm$ 0.0072 0.5277 $\pm$ 0.0067 0.5217 $\pm$ 0.0056
@@
 Dual 0.5192 $\pm$ 0.0092 0.5792 $\pm$ 0.0093 0.5736 $\pm$ 0.0075
+Dual 0.1936 $\pm$ 0.1637 0.2627 $\pm$ 0.1824 0.2858 $\pm$ 0.1467
+Dual 0.2578 $\pm$ 0.0340 0.3363 $\pm$ 0.0270 0.3431 $\pm$ 0.0316
 SEED-V 5-Class Dual 0.2735 $\pm$ 0.0032 0.4235 $\pm$ 0.0022 0.4137 $\pm$ 0.0023
@@
 Dual 0.2561 $\pm$ 0.0051 0.4106 $\pm$ 0.0042 0.4019 $\pm$ 0.0034
+Dual 0.2062 $\pm$ 0.0099 0.3707 $\pm$ 0.0075 0.3620 $\pm$ 0.0083
+Dual 0.2212 $\pm$ 0.0076 0.3826 $\pm$ 0.0057 0.3757 $\pm$ 0.0052
 ISRUC_S3 5-Class Dual 0.7671 $\pm$ 0.0091 0.8202 $\pm$ 0.0071 0.7856 $\pm$ 0.0031
@@
 Dual 0.7359 $\pm$ 0.0324 0.7950 $\pm$ 0.0259 0.7621 $\pm$ 0.0311
+Dual 0.6218 $\pm$ 0.0427 0.6956 $\pm$ 0.0279 0.6664 $\pm$ 0.0316
+Dual 0.5478 $\pm$ 0.0302 0.6429 $\pm$ 0.0307 0.6258 $\pm$ 0.0448
+4.4 Ablation Study
+We conduct ablation studies on three datasets with the same five seeds as in the main experiments to evaluate the key components of CodeBrain (Table 3). Below is the detailed analysis:
+(1) Tokenizer configuration: we compare the proposed TFDual-Tokenizer (Dual) with variants using a single domain codebook (Temporal or Frequency), or a shared codebook jointly reconstructing both domains (Mixed).
+Across all datasets, the Dual codebook consistently yields superior performance.
+(2) Contrastive learning (CL): We evaluate the impact of contrastive learning in TFDual-Tokenizer pretraining.
+It leads to consistent gains, indicating a better capture of temporal patterns.
+Moreover, the CL facilitates convergence of the temporal codebook, with detailed analyses provided in Appendix E.
+(3) Components of EEGSSM: Evaluating the SWA, SGConv, and Gate modules, which demonstrate improvements in EEG representation learning.
+(4) Scaling Laws: Prior works (wang2024cbramod, ; jianglarge, ) explored scaling with 1-1000 hours of EEG data for pretraining.
+We extend to 1k-9k hours and 3M-150M models.
+As shown in Figure 3, Kappa consistently improves with more data and parameters.
+These results confirm that larger models yield consistent but diminishing returns.
+More detailed results are in Appendix M.
+We further investigate several key design choices, including mask ratio, SWA window size, codebook size, and patch size in Appendix J.
+We report robustness experiments to demonstrate the adaptability of dynamic positional embeddings in Appendix L, and computational efficiency in Appendix K.
+We also provide low-resource comparisons (Appendix N) and pretraining curves (Appendix D), showing stable VQ convergence and efficient EEGSSM training.
 Figure 3: Model and training data scaling laws of CodeBrain across three datasets on Cohen’s Kappa.
 4.5 Vector Visualization
+To demonstrate how the TFDual-Tokenizer models heterogeneous EEG, we visualize its learned temporal and frequency codes on the ISRUC_S3 dataset.
+As shown in Figure 4(a)(b), each domain-specific codebook captures meaningful structures: temporal codes correspond to neural events (e.g., slow waves), while frequency codes highlight spectral rhythms such as dominant delta activity, either of which can support sleep stage discrimination.
+Yet in many cases a single domain is insufficient, and discrimination arises only from their composition, as in Figure 4(c)(d), where the same temporal code can pair with different frequency codes, and vice versa, to yield complementary representations.
+This decoupled design expands the representation space and enhances interpretability, with additional potential links to neurophysiological features and quantitative analyses provided in Appendix A.
 Figure 4: Decoupled time-frequency codes visualization on ISRUC_S3 dataset.
-To better understand how the proposed TFDual-Tokenizer overcomes the limitations of single-codebook quantization in handling heterogeneous EEG signals, we visualize its learned temporal and frequency codes on the ISRUC_S3 dataset.
-As shown in Figure 4 (a) and (b), both temporal and frequency codebooks exhibit clustering behavior, capturing domain-specific features, such as characteristic waveforms or frequency patterns associated with specific sleep stages (e.g., N3).
-Figure 4 (c) and (d) further demonstrate that our decoupled design enables flexible composition: the same temporal code can pair with different frequency codes to represent different sleep stages, and vice versa.
-This indicates that temporal and frequency representations contribute both independently and complementarily to the final task representation.
-Such tokenizer supports a richer and more structured representation space for heterogeneous features, improving both performance and interpretability.
 5 Conclusion
-In this work, we propose CodeBrain, an efficient EEG foundation model that integrates three key components: a temporal–frequency decoupled tokenizer for heterogeneous representation learning, a structured state space backbone for modeling sparse global dependencies, and a sliding window attention mechanism for capturing local intra-patch patterns.
-Through comprehensive evaluation, CodeBrain consistently outperforms prior foundation models under linear probing, demonstrating strong generalization and transferability.
-Our results highlight the potential of CodeBrain as a versatile backbone for EEG-based brain–computer interface applications, offering a scalable and interpretable foundation for future developments in neural decoding and AI in neuroscience.
-Appendix A Preliminaries
-Convolution State Space Models
+In this paper, we present CodeBrain, a EEG foundation model that unifies interpretable tokenization with a brain-inspired multi-scale architecture.
+The TFDual-Tokenizer decouples heterogeneous EEG signals, expanding the representation space while suggesting domain-specific interpretability, and the EEGSSM architecture integrates structured global convolution with sliding-window attention to efficiently capture both long-range and local dependencies.
+Pretrained on the large-scale TUEG corpus, CodeBrain demonstrates strong generalization across 8 tasks and 10 datasets under distribution shifts, with comprehensive ablations, scaling-law analyses confirming robustness and scalability.
+These results establish CodeBrain as a strong foundation for neural time-series representation learning.
+Appendix A Interpretability Analyses of TFDual-Tokenizer Representations
+We conduct interpretability analyses of the decoupled TFDual-Tokenizer.
+As case studies, we visualize selected tokens in relation to well-established physiological markers, showing that frequency codes tend to align with spectral rhythms, whereas temporal codes correspond to neural events.
+These qualitative examples highlight that the tokenizer does not discretize signals in an abstract manner, but captures domain-specific structures with clear clinical relevance.
+Finally, we complement these visualizations with a quantitative analysis of class-specific token usage on 4 datasets, demonstrating that the learned tokens are not only interpretable but also discriminative, contributing to improved downstream classification.
+A.1 Exploring Frequency Token Patterns in Relation to Spectral Rhythms
+To illustrate the interpretability of the learned frequency tokens, we take sleep staging as an example downstream task and analyze token embeddings.
+We focus on N2 and N3 stages, since they are characterized by the most distinctive spectral rhythms in clinical sleep scoring: spindle in the sigma band (11-16 Hz) for N2, and slow wave in the delta band (0.5-4 Hz) for N3.
+As shown in Figure 5, frequency code No.
+2298 captures a sigma bump, with prominent peaks localized in the spindle range.
+This aligns with the definition of N2 [68], which identifies spindle as a key marker.
+Quantitatively, this code exhibited a clear sigma peak in approximately $15.4\%$ of its assigned patches.
+Similarly, Figure 6 shows that frequency code No.
+32 predominantly encodes delta dominance, a typical spectral feature of N3 sleep.
+This code displayed clear delta dominance in about $18.3\%$ of its assigned segments.
+Taken together, these findings suggest that our frequency-branch tokenizer helps establish a frequency vocabulary for EEG, suggesting potential links between the learned codes and clinically interpretable oscillations beyond class frequency bias.
+Figure 5: Frequency code capturing sigma bump, a typical spectral rhythm of N2 stage.
+Figure 6: Frequency code capturing delta dominance, a typical spectral rhythm of N3 stage.
+A.2 Exploring Temporal Token Patterns in Relation to Neural Events
+In addition to spectral rhythms, we further analyzed temporal codes on the sleep staging dataset to assess whether the tokenizer captures characteristic neural events in raw EEG.
+Figure 8 and Figure 7 show that certain codes are closely associated with K-complexes and sleep spindles, the hallmark waveforms of N2 sleep.
+Quantitatively, $\approx 5.34\%$ of the assigned patches contained K-complexes and $\approx 7.01\%$ contained spindles.
+These rates are consistent with clinical expectations (1–2 K-complexes and 2–5 spindles per minute in N2 [68]) and, when considering the overall prevalence of N2 in five-class sleep staging, still substantially exceed what would be expected from random token assignment.
+Figure 7: Temporal code capturing sleep spindle, a typical neural events of N2 stage.
+Figure 8: Temporal code capturing K complex, a typical neural events of N2 stage.
+Figure 9 shows that temporal code No.
+1537 corresponds to slow waves, the defining feature of N3 sleep, occurring in $\approx 40.4\%$ of its assigned patches.
+This exceeds the $20\%$ per-epoch criterion for N3 scoring [68] and, given the overall prevalence of N3, indicates that this token provides a meaningful link to neural events.
+Figure 9: Temporal code capturing slow wave, a typical neural events of N3 stage.
+Taken together, these observations suggest that the temporal branch of our tokenizer contributes to establishing a vocabulary of neurophysiological events, complementing the frequency-domain findings.
+While temporal waveforms are often noisier and harder to model than spectral rhythms [16], our decoupled design allows temporal tokens to emerge with meaningful associations to critical neural events.
+This indicates that the temporal branch offers useful links to clinically relevant waveforms and demonstrates the effectiveness of our approach in capturing complementary structure.
+A.3 Class-Specific Token Ratio Analysis for TFDual-Tokenizer
+To better understand the contribution of the TFDual-Tokenizer module, which decouples time and frequency domain representations, in learning discriminative EEG features, we conduct an analysis of class-specific tokens.
+A token (code) is considered class-specific if it predominantly appears in samples of a single class.
+Formally, for a given token $c$, let $N_{c}^{(y)}$ denote the number of times $c$ appears in class $y$.
+The dominance ratio is defined as:
+$\text{Dominance}(c)=\frac{\max_{y}N_{c}^{(y)}}{\sum_{y}N_{c}^{(y)}}$ (17)
+If $\text{Dominance}(c)>\tau$ (we use $\tau=1$), the token is deemed class-specific. The class-specific ratio for a codebook is then computed as:
+$\text{Class-Specific Token Ratio}=\frac{\text{\# class-specific tokens}}{\text{Total tokens in codebook}}$ (18)
+Figure 10 illustrates the proportion of class-specific tokens derived from three configurations: using only the temporal codebook, only the frequency codebook, and a combination of both (TF-Decoupled).
+We employ two independent codebooks to capture complementary information via the proposed TFDual-Tokenizer module.
+Across all four datasets, including two for emotion recognition (FACED and SEED-V) and two for sleep staging (ISRUC_S3 and ISRUC_S1), the decoupled codebook consistently achieves the highest class-specific token ratios, reaching 54.7% on ISRUC_S3 and 46.4% on FACED.
+These results confirm that decoupling temporal and frequency domain information significantly enhances the model’s ability to capture class-discriminative patterns.
+Temporal and frequency domains exhibit inherently heterogeneous characteristics: temporal signals capture transient waveform dynamics, while frequency components reflect oscillatory structures over broader time scales.
+By decoupling these two domains and learning separate token representations, the model is able to preserve their distinct informational content.
+The TFDual-Tokenizer then integrates temporally localized patterns—such as emotion-related bursts or sleep-stage-specific oscillations—that are highly class-discriminative.
+The TFDual-Tokenizer enables richer and more effective EEG representations.
+Figure 10: Class-Specific Code Ratio Across Different Codebooks.
+Appendix B Structure State Space Model
 The state-space model is a classic model in control theory, and it represents the operational state of a system using first-order differential equations (ODE).
@@
 Now, Equation 20 resembles an architecture similar to RNN, allowing us to recurrently compute $x_{k}$.
-Let the initial state be $x_{k-1}=0$, and we can unroll Equation 18 as follows:
+Let the initial state be $x_{k-1}=0$, and we can unroll Eq 20 as follows:
 $y_{k}=\overline{CA^{k}B}u_{0}+\overline{CA^{k-1}B}u_{1}+...+\overline{CAB}u_{k-1}+\overline{CB}u_{k}$ (20)
@@
 Following [31, 70], we hope to parameterize $K$ directly rather than through $\{A,B,C\}$ because we can eliminate complex parameterization and accelerate the entire convolution.
-Appendix B Pretraining Results
+While Eq. 22 shows that the SSM can be computed using FFT-based convolution, it is important to formalize the complexity guarantee. We now state the following proposition.
+Proposition 1.
+Let $(A,B,C,D)$ denote the discretized state-space matrices of an S4 layer, with input sequence $u\in\mathbb{R}^{N}$.
+The output $y\in\mathbb{R}^{N}$ can be written as a linear convolution y=kuy=k*u with kernel
+$k_{0}=D,\quad k_{n}=CA^{n-1}B,\quad n\geq 1.$
+Using the convolution theorem and the FFT, this convolution can be computed in time $\Theta(N\log N)$.
+Proof. This proof follows the proof in Lemma C.2 of gu2021efficiently [49]:
+Expanding the recurrence
+$x_{n+1}=Ax_{n}+Bu_{n},\qquad y_{n}=Cx_{n}+Du_{n},$ (23)
+yields
+$y_{n}=Du_{n}+\sum_{t=0}^{n-1}CA^{\,n-1-t}B\,u_{t}=\sum_{t=0}^{n}k_{n-t}\,u_{t},$ (24)
+where $k_{0}=D$ and $k_{n}=CA^{n-1}B$. Thus y=kuy=k*u.
+Let $\mathcal{F}_{N}$ denote the $N$-point DFT. By the convolution theorem,
+$\mathcal{F}_{N}(y)=\mathcal{F}_{N}(k)\odot\mathcal{F}_{N}(u),$ (25)
+so that
+$y=\mathcal{F}_{N}^{-1}\!\left(\mathcal{F}_{N}(k)\odot\mathcal{F}_{N}(u)\right).$ (26)
+The cost of forward FFT, element-wise product, and inverse FFT is $\Theta(N\log N)$.
+The S4 parameterization ensures $A$ admits a diagonal-plus-low-rank (DPLR) structure, so the kernel takes the form
+$k_{n}=\sum_{j=1}^{r}\alpha_{j}\,\lambda_{j}^{\,n-1},\qquad r\ll N.$ (27)
+Each exponential sequence can be generated recursively in $\mathcal{O}(N)$, yielding all $k_{0},\ldots,k_{N-1}$ in linear time.
+Combining the above,
+$T(N)=\underbrace{\Theta(N\log N)}_{\text{FFT convolution}}+\underbrace{\mathcal{O}(N)}_{\text{kernel generation}}=\Theta(N\log N).$ (28)
+Appendix C Theoretical Analysis of Decoupled Codebook Training
+Proposition 2.1.
+Let $X=(X_{t},X_{f})$ denote the temporal and frequency representations of an EEG segment, assumed approximately independent.
+Consider an additive reconstruction distortion $d(x,\hat{x})=d_{t}(x_{t},\hat{x}_{t})+d_{f}(x_{f},\hat{x}_{f})$.
+Under a fixed total codebook size $K=2^{R}$, a product codebook $\mathcal{C}_{t}\times\mathcal{C}_{f}$ with $|\mathcal{C}_{t}|=2^{R_{t}}$ and $|\mathcal{C}_{f}|=2^{R_{f}}$, $R_{t}+R_{f}=R$, achieves a minimum expected distortion
+$D_{\text{prod}}^{\star}(R)=\min_{R_{t}+R_{f}=R}\Bigl(D_{t}^{\star}(R_{t})+D_{f}^{\star}(R_{f})\Bigr),$ (29)
+which satisfies
+$D_{\text{mix}}^{\star}(R)\;\geq\;D_{\text{prod}}^{\star}(R),$ (30)
+where $D_{\text{mix}}^{\star}(R)$ is the minimum distortion of a single mixed codebook $\mathcal{C}_{\text{mix}}\subset\mathbb{R}^{d_{t}+d_{f}}$ of size $2^{R}$.
+Proof.
+The argument combines the separability of the rate–distortion (R–D) function for independent sources under additive distortion and the high-rate quantization approximation to the Shannon R–D limit.
+Let $X=(X_{t},X_{f})$ with $X_{t}\perp X_{f}$, and distortion
+$d(x,\hat{x})\;=\;d_{t}(x_{t},\hat{x}_{t})+d_{f}(x_{f},\hat{x}_{f}).$ (31)
+Then the Shannon R–D function satisfies
+$R(D)\;=\;\min_{D_{t}+D_{f}=D}\,\Bigl(R_{t}(D_{t})+R_{f}(D_{f})\Bigr),$ (32)
+where $R_{t}(\cdot)$ and $R_{f}(\cdot)$ are the marginal R–D functions for $X_{t}$ and $X_{f}$. By convex duality, the optimal test channel factorizes:
+$p(\hat{x}_{t},\hat{x}_{f}\mid x_{t},x_{f})\;=\;p(\hat{x}_{t}\mid x_{t})\,p(\hat{x}_{f}\mid x_{f}),$ (33)
+and the optimal distortion allocation solves
+$\min_{D_{t},D_{f}}\;R_{t}(D_{t})+R_{f}(D_{f})\quad\text{s.t. }D_{t}+D_{f}=D,$ (34)
+with KKT condition $R_{t}^{\prime}(D_{t}^{\star})=R_{f}^{\prime}(D_{f}^{\star})$.
+At rate $R_{t}$ bits for $X_{t}$ and $R_{f}$ bits for $X_{f}$ (with $R_{t}+R_{f}=R$), the minimal achievable distortions are $D_{t}^{\star}(R_{t})$ and $D_{f}^{\star}(R_{f})$.
+In the high-rate regime, practical vector quantizers approach these Shannon limits, yielding
+$D_{\text{prod}}(R_{t},R_{f})\;\approx\;D_{t}^{\star}(R_{t})+D_{f}^{\star}(R_{f}).$ (35)
+Optimizing over all feasible splits gives
+$D_{\text{prod}}^{\star}(R)\;=\;\min_{R_{t}+R_{f}=R}\bigl(D_{t}^{\star}(R_{t})+D_{f}^{\star}(R_{f})\bigr).$ (36)
+Any mixed codebook $\mathcal{C}_{\text{mix}}\subset\mathbb{R}^{d_{t}+d_{f}}$ with $|\mathcal{C}_{\text{mix}}|=2^{R}$ cannot beat the Shannon R–D limit, so
+$D_{\text{mix}}^{\star}(R)\;\geq\;D^{\star}(R).$ (37)
+But from (34)– (35),
+$D^{\star}(R)\;=\;\min_{R_{t}+R_{f}=R}\bigl(D_{t}^{\star}(R_{t})+D_{f}^{\star}(R_{f})\bigr)\;=\;D_{\text{prod}}^{\star}(R).$ (38)
+Therefore,
+$D_{\text{mix}}^{\star}(R)\;\geq\;D_{\text{prod}}^{\star}(R),$ (39)
+Appendix D Pretraining Results
 Our model follows a two-stage pretraining framework.
@@
 This section reports the pretraining results of both stages, including loss convergence, reconstruction dynamics, and codebook utilization patterns.
-B.1 TFDual-Tokenizer Pretraining Results
+D.1 TFDual-Tokenizer Pretraining Results
 Total Training Loss
@@
 As shown in Figure 13, the frequency codebook demonstrates a rapid decrease in unused codes, while the temporal codebook shows a slower and more incremental reduction.
-A more detailed analysis of the temporal–frequency complementarity is provided in Section C.
+A more detailed analysis of the temporal-frequency complementarity is provided in Section A.3.
 Figure 13: Unused code dynamics of the TFDual-Tokenizer.
-B.2 EEGSSSM Pretraining Results
+D.2 EEGSSSM Pretraining Results
 We plot the pretraining loss curve of EEGSSM in Figure 14.
@@
 Figure 14: Pretraining Loss Curve of EEGSSM.
-Appendix C Class-Specific Token Ratio Analysis for TFDual-Tokenizer
-To better understand the contribution of the TFDual-Tokenizer module, which decouples time and frequency domain representations, in learning discriminative EEG features, we conduct an analysis of class-specific tokens.
-A token (code) is considered class-specific if it predominantly appears in samples of a single class.
-Formally, for a given token $c$, let $N_{c}^{(y)}$ denote the number of times $c$ appears in class $y$.
-The dominance ratio is defined as:
-$\text{Dominance}(c)=\frac{\max_{y}N_{c}^{(y)}}{\sum_{y}N_{c}^{(y)}}$ (21)
-If $\text{Dominance}(c)>\tau$ (we use $\tau=1$), the token is deemed class-specific. The class-specific ratio for a codebook is then computed as:
-[…]
-Figure 9 illustrates the proportion of class-specific tokens derived from three configurations: using only the temporal codebook, only the frequency codebook, and a combination of both (TF-Decoupled).
-We employ two independent codebooks to capture complementary information via the proposed TFDual-Tokenizer module.
-Across all four datasets, including two for emotion recognition (FACED and SEED-V) and two for sleep staging (ISRUC_S3 and ISRUC_S1), the decoupled codebook consistently achieves the highest class-specific token ratios, reaching 54.7% on ISRUC_S3 and 46.4% on FACED.
-These results confirm that decoupling temporal and frequency domain information significantly enhances the model’s ability to capture class-discriminative patterns.
-Temporal and frequency domains exhibit inherently heterogeneous characteristics: temporal signals capture transient waveform dynamics, while frequency components reflect oscillatory structures over broader time scales.
-By decoupling these two domains and learning separate token representations, the model is able to preserve their distinct informational content.
-The TFDual-Tokenizer then integrates temporally localized patterns—such as emotion-related bursts or sleep-stage-specific oscillations—that are highly class-discriminative.
-The TFDual-Tokenizer enables richer and more effective EEG representations.
-Figure 9: Class-Specific Code Ratio Across Different Codebooks.
-Appendix D Improving Temporal Codebook Learning via Contrastive Loss
+Appendix E Improving Temporal Codebook Learning via Contrastive Loss
 To improve the learning of the temporal codebook in our TFDual-Tokenizer, we introduce a contrastive loss as one of the objectives during pretraining.
@@
 Figure 15: Effect of Contrastive Loss on Temporal Codebook Learning.
-Appendix E Dataset Description
+Appendix F Dataset Description
 We evaluate the CodeBrain across eight diverse downstream tasks covering ten publicly available EEG datasets.
@@
 We follow the official train-test split and further divide the training set into training and validation subsets using an 8:2 subject-wise ratio, consistent with existing benchmarks.
-Appendix F Baselines and Metrics Description
-F.1 Metrics
+Appendix G Baselines and Metrics Description
+G.1 Metrics
 To comprehensively evaluate our model, we compare it with a set of strong baselines commonly used in EEG analysis.
@@
 For model selection and comparison, AUROC is used as the main evaluation metric for binary classification tasks, and Cohen’s Kappa is used for multi-class scenarios.
-F.2 Baselines
+G.2 Baselines
 We compare our CodeBrain model against 5 publicly available EFMs that have released pre-trained weights, covering a diverse set of pretraining strategies.
@@
 It introduces a criss-cross transformer architecture comprising parallel spatial and temporal attention mechanisms, enabling separate yet simultaneous modeling of spatial and temporal relationships.
-Appendix G Hyperparameter Setting
+Appendix H Hyperparameter Setting
 We provide detailed hyperparameter configurations for the two-stage pretraining of our CodeBrain model and the fine-tuning settings across ten downstream tasks.
-G.1 Pretraining Settings
+H.1 Pretraining Settings
 The pretraining process consists of two stages:
@@
   2. 2.
-    training the EEGSSM framework
+    Training the EEGSSM
 The hyperparameters used in each stage are summarized in Table 4 and Table 5, respectively.
@@
 Data stride 200
-Table 5: Hyperparameters for EEGSSM.
+Table 5: Hyperparameters of Pre-training.
 Hyperparameters Values
@@
 Clipping gradient norm 5
-G.2 Fine-tuning Settings on Downstream Tasks
+H.2 Parameters of EEGSSM
+The model architecture parameters used in EEGSSM during pre-training are shown in Table 6.
+Table 6: Configuration of EEGSSM
+Parameters Values
+Input size 200
+Hidden dimension 200
+Output size 200
+Number of layers 8
+Max sequence length 570
+SGConv state 64
+SGConv bidirectional True
+Layer normalization True
+Sliding window attention length 1s
+H.3 Fine-tuning Settings on Downstream Tasks
 The CodeBrain model is fine-tuned on ten downstream EEG classification tasks, each with task-specific hyperparameters.
@@
 TUAB 1e-5 5e-5 0.4 512
-Appendix H Model Comparison
-Backbone Efficiency Comparison
-To evaluate the computational efficiency of our proposed SGConv module, we conduct an ablation study by replacing it with three common sequence modeling modules: CNN, LSTM, and Transformer.
-We compare their model sizes, floating-point operations (FLOPs), and iteration times, as shown in Figure 11.
-Specifically, the CNN variant uses a three-layer depthwise separable convolution block, while the LSTM and Transformer variants use a single layer of standard LSTM and Transformer Encoder (implemented by Pytorch), respectively.
-In terms of parameter count, SGConv contains only 8.74M parameters, significantly fewer than CNN (28.22M), LSTM (30.35M), and Transformer (27.79M).
-For FLOPs, SGConv also achieves the lowest computational cost at 15.17G, compared to Transformer’s 21.2G.
-Regarding iteration time, SGConv is slightly slower than Transformer and CNN models in terms of training speed, but it outperforms the LSTM model.
-In summary, SGConv effectively reduces the number of parameters while maintaining computational complexity, which helps the model to be trained and inferred on smaller GPUs.
-Figure 11: Computational Overhead of Using Different Backbones in the EEGSSM Module.
-Additional Evaluation on Other BCI Tasks
-We report the performance of CodeBrain on four additional EEG datasets not included in the main text, covering diverse domains of sleep staging, motor imagery, event detection, and abnormality classification in Tables 7 to 10.
+Appendix I Additional Evaluation on Other BCI Tasks
+We report the performance of CodeBrain on four additional EEG datasets not included in the main text, covering diverse domains of sleep staging, motor imagery, event detection, and abnormality classification in Tables 8 to 11.
 These allow us to assess the cross-domain generalization ability of our pretrained model beyond the main text.
 We note that both TUAB and TUEV originate from the TUH EEG corpus [33], which overlaps with our pretraining source (TUEG).
-To avoid overfitting to this distribution and promote generalization, we stop pretraining at epoch 10 as discussed in Section B.2.
+To avoid overfitting to this distribution and promote generalization, we stop pretraining at epoch 10 as discussed in Section D.2.
 While this may limit gains on TUH datasets compared to previous EFM, such as CBraMod (trained for 40 epochs in the same pretraining dataset) [15], CodeBrain still achieves superior or competitive results.
@@
 Although its Balanced Accuracy is lower than CBraMod [15], we attribute this to reduced sensitivity on the rare SPSW class.
-Since TUEV shares distributional overlap with our pretraining source (TUEG), we stop pretraining at epoch 10 to prevent overfitting, unlike CBraMod’s 40-epoch training as discussed in subsection B.2.
+Since TUEV shares distributional overlap with our pretraining source (TUEG), we stop pretraining at epoch 10 to prevent overfitting, unlike CBraMod’s 40-epoch training as discussed in subsection D.2.
 We also report LaBraM’s results based on its original 23-channel setting [16], while CodeBrain follows the 16-channel configuration used in CBraMod.
@@
 While such architectural choices may enhance performance, we follow their respective fine-tuning settings.
+In addition, following the experimental setup of CBraMod, we also report the results of experiments conducted by removing all TUEV and TUAB samples from the TUEG dataset.
+It can be seen that although our model’s performance slightly declined on TUEV, it still surpassed other baselines.
 In such cases, CodeBrain still outperforming both LaBraM and EEGPT under their own fine-tuning settings clearly demonstrates the robustness of our approach.
@@
 CBraMod[15] 0.6772 ± 0.0096 0.8342 ± 0.0064 0.6671 ± 0.0107
+CodeBrain (Excluding) 0.6838 ± 0.0291 0.8293 ± 0.0163 0.6375 ± 0.0182
 CodeBrain 0.6912 ± 0.0101 0.8362 ± 0.0048 0.6428 ± 0.0062
@@
 Similar to TUEV, TUAB is part of the TUH EEG corpus family [33] and thus closely aligned with our pretraining source (TUEG).
-As discussed in subsection B.2, we adopt an early stopping strategy at epoch 10 to mitigate overfitting to this distribution, which may partly account for the slightly lower AUROC and AUC-PR compared to CBraMod, trained for 40 epochs on the same dataset.
+As discussed in subsection D.2, we adopt an early stopping strategy at epoch 10 to mitigate overfitting to this distribution, which may partly account for the slightly lower AUROC and AUC-PR compared to CBraMod, trained for 40 epochs on the same dataset.
 While LaBraM leverages a 23-channel montage [16] and EEGPT [14] employs two convolutional layers before the foundation model, we retain their respective fine-tuning protocols for comparison.
+In addition, the impact of duplicate data in the pre-training set on our model is smaller on the TUAB dataset, and in some experiments it can even surpass situations without leaking.
+This may be due to the larger size of the TUAB dataset.
 Despite these potentially stronger configurations, CodeBrain still exceeds both models under their own settings, highlighting its strong and consistent generalization.
@@
 CBraMod[15] 0.8289 ± 0.0022 0.9258 ± 0.0008 0.9227 ± 0.0011
+CodeBrain (Excluding) 0.8288 ± 0.0064 0.9061 ± 0.0039 0.9012 ± 0.0020
 CodeBrain 0.8294 ± 0.0013 0.9100 ± 0.0006 0.9030 ± 0.0009
-Appendix I Ablation on Mask Ratio
+Appendix J Ablation on Design Choices
+J.1 Ablation on Mask Ratio
 We conduct an ablation study to investigate the effect of the mask ratio in the EEGSSM pretraining framework.
@@
 Mask Ratio Cohen’s Kappa Weighted F1 Balanced Accuracy
-0.1 0.5056 ± 0.0054 0.5652 ± 0.0063 0.5579 ± 0.0059
-0.2 0.5111 ± 0.0042 0.5721 ± 0.0030 0.5692 ± 0.0035
-0.3 0.5196 ± 0.0081 0.5757 ± 0.0059 0.5707 ± 0.0072
-0.4 0.5253 ± 0.0042 0.5831 ± 0.0039 0.5795 ± 0.0046
-0.5 0.5270 ± 0.0039 0.5849 ± 0.0033 0.5816 ± 0.0028
-0.6 0.5156 ± 0.0066 0.5720 ± 0.0075 0.5671 ± 0.0095
-0.7 0.5113 ± 0.0087 0.5699 ± 0.0063 0.5621 ± 0.0075
-0.8 0.5026 ± 0.0057 0.5467 ± 0.0044 0.5408 ± 0.0042
-0.9 0.4906 ± 0.0083 0.5369 ± 0.0084 0.5345 ± 0.0095
+0.1 0.5184 $\pm$ 0.0039 0.5746 $\pm$ 0.0158 0.5692 $\pm$ 0.0064
+0.2 0.5239 $\pm$ 0.0036 0.5834 $\pm$ 0.0071 0.5821 $\pm$ 0.0040
+0.3 0.5327 $\pm$ 0.0100 0.5859 $\pm$ 0.0036 0.5836 $\pm$ 0.0076
+0.4 0.5391 $\pm$ 0.0045 0.5938 $\pm$ 0.0057 0.5904 $\pm$ 0.0058
+0.5 0.5406 $\pm$ 0.0084 0.5953 $\pm$ 0.0113 0.5941 $\pm$ 0.0098
+0.6 0.5295 $\pm$ 0.0075 0.5822 $\pm$ 0.0090 0.5793 $\pm$ 0.0112
+0.7 0.5242 $\pm$ 0.0077 0.5800 $\pm$ 0.0065 0.5744 $\pm$ 0.0096
+0.8 0.5157 $\pm$ 0.0065 0.5564 $\pm$ 0.0067 0.5528 $\pm$ 0.0040
+0.9 0.5034 $\pm$ 0.0078 0.5457 $\pm$ 0.0084 0.5451 $\pm$ 0.0114
 Table 13: Performance of CodeBrain on SEED-V Dataset under Different Mask Ratios.
@@
 Figure 16: Performance Across Different Mask Ratios on FACED, SEED-V, and ISRUC_S3.
-Figure 13: EEGSSM Pre-Training Loss Curve for Different Mask Ratios.
 To further illustrate this pattern, we visualize the training loss curves across different mask ratios in Figure 17.
@@
 Moderate masking strikes a balance by being sufficiently challenging to promote abstraction, while still being learnable, thereby facilitating better generalization across downstream tasks.
-Appendix J Detailed Results on Scaling Laws
+Figure 17: EEGSSM Pre-Training Loss Curve for Different Mask Ratios.
+J.2 Ablation on SWA Window Size
+We conduct an ablation study to investigate the effect of the SWA window size in the EEGSSM framework.
+The window size of SWA means the length of the segment observed by the attention mechanism in the SWA mechanism.
+Usually, the window size of SWA is an odd number because the model generally focuses not only on a single segment but also needs to observe features in adjacent segments.
+Therefore, its size is usually $2n+1$, where $n$ represents the length of SWA.
+As shown in Figure 18 and Figure 19, SWA window size = 1 achieved the best performance on both datasets among the 5 SWA window sizes.
+With the increase of the SWA window size, the SWA is close to self-attention.
+Therefore, as the SWA window size increases, the model’s performance actually improves because more data will be involved in the attention calculation.
+When the SWA window size is 1, it is a special case where SWA only calculates for one second.
+This is equivalent to the model performing attention mechanism calculations within one second.
+Overall, SWA can help the model achieve a certain performance improvement.
+Figure 18: Performance of different SWA window sizes on the FACED dataset.
+Figure 19: Performance of different SWA window sizes on the SEED-V dataset.
+J.3 Ablation on Codebook Size
+The size of the codebook is an important parameter; a codebook that is too large may lead to unstable training, while a codebook size that is too small may result in mixed information.
+Ideally, the codebook should maintain a small amount of unused codes but not be 0.
+We tested several combinations of different time-domain and frequency-domain codebook sizes to observe their unused codes during Tokenizer training.
+Table 15 shows the unused temporal codes and unused frequency codes under different codebook size combinations.
+When both the Temporal codebook size and Frequency codebook size are set to 2048, the unused frequency code is 0, indicating that there are duplicate frequency codes in the current codebook.
+When choosing a codebook size of 8192, the unused temporal codes reached 3401 and the unused frequency codes reached 1260, indicating that a large number of codes in the completed training codebook were not used.
+We ultimately selected 4096 for both.
+Although the frequency domain often yields richer representations, enlarging its codebook may increase reliance on it, so we kept temporal and frequency codebooks equal.
+Considering capacity and utilization, 4096–4096 offers the best trade-off.
+Table 15: Temporal and frequency codebook statistics
+Temporal Codebook Size Frequency Codebook Size Unused Temporal Codes Unused Frequency Codes
+2048 2048 12 0
+2048 4096 116 0
+4096 4096 225 165
+4096 8192 321 931
+8192 8192 3401 1260
+J.4 Ablation on Patch Size
+The size of the patch window is also an important adjustable parameter, which affects temporal resolutions and masking strategies.
+To explore the impact of patch window size on the model, we used window sizes ranging from 0.5s to 5s for complete two-stage pre-training and full-parameter fine-tuning.
+Note that for some datasets where the patch size is larger than the channel length, such as SEED-V, we pad the portion exceeding the available data to match the patch size in this experiment.
+Table 16 and Table 17 show the performance of our method with different patch sizes on the SEED-V dataset and ISRUC_3 dataset.
+Notably, in SEED‑V, patch lengths longer than 1s require heavy padding, causing large performance drops; in ISRUC_S3, the shortest 0.5s patches achieve the worst performance, likely because they fragment key waveforms in sleep staging (e.g., Spindle $\geq$0.5s).
+From these results, the 1s setting is supported by two key considerations:
+Broad compatibility with downstream task.
+1s is a divisor of most downstream sequence lengths (1–30s), minimizing padding and ensuring transferability.
+For example, on the SEED-V dataset, if the patch size chosen by the model is greater than 1s, some methods (such as padding) need to be adopted to enable model training.
+These methods may usually impair the model’s performance because they introduce additional noise or increase computational load.
+EEG datasets with durations less than 1 second are relatively rare, as most datasets have at least 1 second of data.
+If the data does not exactly match the whole seconds, the cost of processing such a dataset is also relatively small.
+Prior EEG foundation models (e.g., LaBram [16], CBraMod [15]) also adopt 1s patches for this reason.
+Semantic integrity.
+Choosing a 1s patch length preserves the natural structure of EEG waveforms and prevents semantic fragmentation.
+Many physiologically meaningful EEG events have characteristic durations: for example, K-complexes are typically around 1s, spindles last 0.5–2s, and event-related potentials such as P300 occur in the range of 0.3–0.6s.
+If patches are shorter than these characteristic scales, the temporal branch of the TFDual-Tokenizer may only capture partial fragments of these waveforms, leading to loss of semantic context.
+Table 16: Performance on SEED-V with different patch sizes
+Patch Length Cohen’s Kappa Weighted F1 Balanced Accuracy
+[…]
+[…]
+[…]
+[…]
+Table 17: Performance on ISRUC_3 with different patch sizes
+Patch Length Cohen’s Kappa Weighted F1 Balanced Accuracy
+[…]
+[…]
+[…]
+[…]
+Appendix K Backbone Efficiency Comparison
+To evaluate the computational efficiency of our proposed SGConv module, we conduct an ablation study by replacing it with three common sequence modeling modules: CNN, LSTM, and Transformer in EEGSSM block.
+We compare their model sizes, floating-point operations (FLOPs), and iteration times, as shown in Figure 20.
+Specifically, the CNN variant uses a 3-layer depthwise separable convolution block, while the LSTM and Transformer variants use a single layer of standard LSTM and Transformer Encoder (implemented by Pytorch), respectively.
+In terms of parameter count, SGConv contains 15.17M parameters, fewer than CNN (16.22M), LSTM (17.35M), and Transformer (21.2M).
+For FLOPs, SGConv also achieves the lowest computational cost at 8.74G, compared to Transformer’s 27.79G.
+Regarding iteration time, SGConv is slightly slower than Transformer and CNN models in terms of training speed, but it outperforms the LSTM model.
+In summary, SGConv effectively reduces the number of parameters while maintaining computational complexity, which helps the model to be trained and inferred on smaller GPUs.
+Figure 20: Computational Overhead of Using Different Backbones in the EEGSSM Module.
+Appendix L Channel Robustness
+In real-world scenarios, the collection of EEG often encounters situations where channels are missing, especially when using machines from different manufacturers.
+To test the model’s performance on datasets with missing channel data, we randomly mask some channels in the training data for full parameter fine-tuning.
+We selected the FACED and SEED-V datasets for experiments because they represent short-sequence and long-sequence cases respectively, and their numbers of channels are relatively complete.
+Figure 21: Performance after randomly masking different numbers of channels during the full parameter fine-tuning stage.
+We evaluate the performance of the CodeBrain model and CBraMod model in scenarios with missing channels.
+We conducted three different experiments, randomly masking 12.5%, 25%, and 50% of the channels in each experiment, respectively.
+Figure 21 shows results of the experiment.
+It can be seen that our CodeBrain model outperforms the CBraMod model in all channel masking scenarios.
+On the FACED dataset, our model’s performance after masking 25% of the channels is still close to that of CBraMod without masking.
+The performance decline of the CBraMod model is also faster than that of our model.
+This trend is even more pronounced on the SEED-V dataset.
+Through this experiment, we can demonstrate that the channel robustness of CodeBrain is stronger than that of the CBraMod model, retaining most of its performance even in cases of channel failure.
+Appendix M Detailed Results on Scaling Laws
 We provide the detailed scaling law results for both data and model size across three representative EEG datasets (FACED, SEED-V, and ISRUC_S3), covering three evaluation metrics.
 For brevity, only Cohen’s kappa scores are included in the main text, while full results are provided in this section.
-Prior work [19, 20] has explored the effect of scaling EEG foundation models using 1 to 1000 hours of pretraining data and model sizes up to 5M parameters.
+Prior work [15, 16] has explored the effect of scaling EEG foundation models using 1 to 1000 hours of pretraining data.
 We extend this analysis in two key dimensions:
@@
   2. 2.
-    Investigating model scaling by varying the depth of the EEGSSM encoder from 3 layers (6.82M parameters) to 8 layers (15.17M parameters).
-J.1 Scaling Laws with Respect to Training Data Volume
+    Investigating model scaling by varying the depth of the EEGSSM encoder from 3 layers (3.86M parameters) to 24 layers (146.75M parameters) and the hidden size from 128 to 384.
+M.1 Scaling Laws with Respect to Training Data Volume
 We examine how the volume of pretraining data influences the downstream performance of CodeBrain.
 Specifically, we scale the pretraining duration from 1k to 9k hours and evaluate the resulting models on three downstream datasets: FACED, SEED-V, and ISRUC_S3.
-Detailed quantitative results across three evaluation metrics (Cohen’s kappa, weighted F1 score, and balanced accuracy) are presented in Tables 1415, and 16, respectively.
+Detailed quantitative results across three evaluation metrics (Cohen’s kappa, weighted F1 score, and balanced accuracy) are presented in Tables 18, 19, and 20, respectively.
 As shown in Figure 22, increasing the amount of pretraining data generally leads to consistent improvements across all datasets and metrics.
@@
 Training Data Cohen’s Kappa Weighted F1 Balanced Accuracy
-1000 Hours 0.4883 ± 0.0102 0.5379 ± 0.0159 0.5344 ± 0.0150
-2000 Hours 0.5001 ± 0.0084 0.5459 ± 0.0107 0.5412 ± 0.0134
-3000 Hours 0.5058 ± 0.0075 0.5606 ± 0.0092 0.5574 ± 0.0086
-4000 Hours 0.5084 ± 0.0018 0.5632 ± 0.0040 0.5600 ± 0.0034
-5000 Hours 0.5034 ± 0.0069 0.5581 ± 0.0114 0.5549 ± 0.0087
-6000 Hours 0.5136 ± 0.0105 0.5682 ± 0.0057 0.5650 ± 0.0071
-7000 Hours 0.5196 ± 0.0114 0.5716 ± 0.0095 0.5694 ± 0.0103
-8000 Hours 0.5201 ± 0.0087 0.5751 ± 0.0065 0.5719 ± 0.0077
-9000 Hours 0.5270 ± 0.0039 0.5849 ± 0.0033 0.5816 ± 0.0028
+1000 Hours 0.5014 $\pm$ 0.0107 0.5462 $\pm$ 0.0146 0.5452 $\pm$ 0.0163
+2000 Hours 0.5133 $\pm$ 0.0068 0.5540 $\pm$ 0.0120 0.5521 $\pm$ 0.0133
+3000 Hours 0.5189 $\pm$ 0.0086 0.5688 $\pm$ 0.0104 0.5687 $\pm$ 0.0083
+4000 Hours 0.5208 $\pm$ 0.0032 0.5741 $\pm$ 0.0041 0.5713 $\pm$ 0.0024
+5000 Hours 0.5171 $\pm$ 0.0040 0.5692 $\pm$ 0.0131 0.5661 $\pm$ 0.0105
+6000 Hours 0.5273 $\pm$ 0.0085 0.5803 $\pm$ 0.0064 0.5764 $\pm$ 0.0069
+7000 Hours 0.5328 $\pm$ 0.0116 0.5854 $\pm$ 0.0082 0.5809 $\pm$ 0.0097
+8000 Hours 0.5336 $\pm$ 0.0082 0.5875 $\pm$ 0.0071 0.5835 $\pm$ 0.0072
+9000 Hours 0.5406 $\pm$ 0.0084 0.5953 $\pm$ 0.0113 0.5941 $\pm$ 0.0098
 Table 19: Training Data Scaling Laws of CodeBrain on SEED-V Dataset.
@@
 Figure 23: EEGSSM Pre-Training Loss Curve for Different Training Data Volume.
-J.2 Scaling Laws with Respect to Model Size
+M.2 Scaling Laws with Respect to Model Size
 We further investigate how model parameters affect downstream performance by scaling the number of layers in the EEGSSM encoder from 3 to 8, resulting in parameter counts ranging from 6.82M to 15.17M.
-Detailed results across the FACED, SEED-V, and ISRUC_S3 datasets are provided in Tables 1718, and 19, respectively.
+Detailed results across the FACED, SEED-V, and ISRUC_S3 datasets are provided in Tables 21, 22, and 23, respectively.
 To visualize the trend more clearly, Figure 24 presents the performance curves as model size increases.
@@
 Table 21: Model Size Scaling Laws of CodeBrain on the FACED Dataset.
-Model Size Cohen’s Kappa Weighted F1 Balanced Accuracy
-6.82M 0.4818 ± 0.0165 0.5362 ± 0.0113 0.5317 ± 0.0182
-8.49M 0.4988 ± 0.0082 0.5506 ± 0.0091 0.5497 ± 0.0084
-10.16M 0.5096 ± 0.0049 0.5686 ± 0.0104 0.5642 ± 0.0067
-11.83M 0.5144 ± 0.0113 0.5705 ± 0.0085 0.5678 ± 0.0095
-13.50M 0.5178 ± 0.0069 0.5772 ± 0.0080 0.5746 ± 0.0067
-15.17M 0.5270 ± 0.0039 0.5849 ± 0.0033 0.5816 ± 0.0028
+Layer Hidden size Params Cohen’s Kappa Weighted F1 Balanced Accuracy
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
 Table 22: Model Size Scaling Laws of CodeBrain on the SEED-V Dataset.
-Model Size Cohen’s Kappa Weighted F1 Balanced Accuracy
-6.82M 0.2609 ± 0.0078 0.4004 ± 0.0112 0.3956 ± 0.0098
-8.49M 0.2638 ± 0.0080 0.4108 ± 0.0091 0.4030 ± 0.0121
-10.16M 0.2645 ± 0.0102 0.4127 ± 0.0158 0.4013 ± 0.0084
-11.83M 0.2663 ± 0.0056 0.4202 ± 0.0068 0.4079 ± 0.0055
-13.50M 0.2724 ± 0.0057 0.4211 ± 0.0060 0.4120 ± 0.0086
-15.17M 0.2735 ± 0.0032 0.4235 ± 0.0022 0.4137 ± 0.0023
+Layer Hidden size Params Cohen’s Kappa Weighted F1 Balanced Accuracy
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
 Table 23: Model Size Scaling Laws of CodeBrain on the ISRUC_S3 Dataset.
-Model Size Cohen’s Kappa Weighted F1 Balanced Accuracy
-6.82M 0.7456 ± 0.0083 0.7862 ± 0.0104 0.7514 ± 0.0084
-8.49M 0.7486 ± 0.0014 0.7942 ± 0.0040 0.7604 ± 0.0035
-10.16M 0.7516 ± 0.0025 0.7985 ± 0.0056 0.7639 ± 0.0079
-11.83M 0.7570 ± 0.0082 0.8064 ± 0.0068 0.7734 ± 0.0061
-13.50M 0.7620 ± 0.0045 0.8153 ± 0.0091 0.7824 ± 0.0100
-15.17M 0.7671 ± 0.0091 0.8202 ± 0.0071 0.7856 ± 0.0031
+Layer Hidden size Params(M) Cohen’s Kappa Weighted F1 Balanced Accuracy
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
 Figure 24: Model Size Scaling Laws on FACED, SEED-V, and ISRUC_S3.
@@
 The loss reduction is particularly evident when increasing from 3 to 6 layers, while the gain starts to saturate beyond 7 layers.
+Even when increased to 24 layers, the reduction in pre-training loss brought by more than 100M parameters is not significant.
 This trend mirrors the downstream performance in Figure 24 and Tables 2123, suggesting that both optimization efficiency and generalization benefit from increased model size—though with diminishing returns as parameter count grows.
@@
 Our results demonstrate consistent improvements in downstream performance as both data volume and model capacity increase, suggesting that EEG foundation models may continue to benefit from further scaling, similar to trends observed in vision and language domains.
-Appendix K Low-Resource Comparison With Existing Methods
+Appendix N Low-Resource Comparison With Existing Methods
 To evaluate the effectiveness of our model under data-scarce conditions, we conduct experiments where only 30% of the training data is available.
-Tables 20 and 21 present the results on the FACED (9-class) and SEED-V (5-class) datasets, respectively, comparing our CodeBrain model with several strong baselines.
+Tables 24 and 25 present the results on the FACED (9-class) and SEED-V (5-class) datasets, respectively, comparing our CodeBrain model with several strong baselines.
 We observe that CodeBrain consistently achieves the best performance among all methods under the 30% setting across all metrics.
@@
 CodeBrain (30%) 0.3356 ± 0.0253 0.4114 ± 0.0225 0.4104 ± 0.0281
-CodeBrain (full) 0.5270 ± 0.0039 0.5849 ± 0.0033 0.5816 ± 0.0028
+CodeBrain (full) 0.5406 $\pm$ 0.0084 0.5953 $\pm$ 0.0113 0.5941 $\pm$ 0.0098
 Table 25: Comparison under Partial Data Setting (30%) on the SEED-V Dataset (5-class).
@@
 CodeBrain (full) 0.2735 ± 0.0032 0.4235 ± 0.0022 0.4137 ± 0.0023
-Appendix L Limitations and Future Work
-Our work presents promising results but also highlights several limitations that offer directions for future exploration.
-(1) More diverse dataset: Although we pre-train on TUEG, the largest publicly available EEG corpus (over 9000 hours) [66], it has limited channel coverage and originates from a single clinical center, leading to potential distributional bias.
-This highlights the need for a more diverse, multi-center pre-training dataset with broader channel and task coverage to support robust cross-domain generalization in EEG research.
-(2) Model scalability: Due to limited data diversity and scale, we used a relatively shallow architecture.
-While our model design supports deeper networks, effectively training them requires significantly larger and more diverse data.
-Given the challenges in acquiring large-scale EEG data, future work may involve collaborative efforts to curate richer pre-training corpora that can support scaling to more expressive EEG foundation models.
-(3) Unimodal focus: Our current framework focuses solely on EEG signals, while other physiological modalities (e.g., EMG, EOG) may also benefit downstream tasks such as emotion recognition and sleep staging.
-Future work will explore incorporating multimodal EXG signals for more comprehensive representation learning for brain foundation models.
-Appendix M Broader impacts
-The development of EEG foundation models such as ours marks an important step forward in the intersection of AI and neurophysiological signal processing.
-Below we outline several potential broader impacts of CodeBrain:
-Novel insights for EEG foundation model.
-While most existing EEG foundation models rely on Transformer-based architectures, our work demonstrates that non-Transformer designs, particularly those inspired by structured convolutional or state-space models, can achieve competitive performance with lower computational overhead.
-This may encourage the community to rethink the current architectural paradigm and stimulate research into inductive biases that are better aligned with EEG signal characteristics.
-We believe this shift could lead to more efficient, interpretable, and scalable EEG modeling frameworks.
-Improving interpretability for clinical trust and scientific insight.
-Our CodeBrain model enhances interpretability through domain-decoupled tokenization, allowing temporal and frequency components of EEG signals to be disentangled and visualized independently.
-This design enables a transparent representation of domain-specific brain features (e.g., delta waves in N3 sleep), supporting both clinical trust and neuroscientific interpretation.
-Such interpretability is critical in high-stakes domains like healthcare, where transparent model decisions facilitate responsible deployment and cross-disciplinary collaboration.
-High availability in the real world.
-By leveraging large-scale pre-training on extensive EEG data, our model can learn general representations applicable to various EEG downstream tasks.
-This capability is crucial for building BCI systems that can be deployed in different clinical settings, thereby enhancing their accessibility and practicality.
-Additionally, through improvements in the model architecture, our model achieves acceptable performance while reducing the parameters of the pre-trained model.
-By decreasing the number of parameters, we lower the memory requirements for deploying the model.